### v1

# Pricing Options on Forwards in Function-Valued Affine Stochastic Volatility Models

Abstract.
We study the pricing of European-style options written on forward contracts within function-valued infinite-dimensional affine stochastic volatility models. The dynamics of the underlying forward price curves are modeled within the Heath-Jarrow-Morton-Musiela framework as solution to a stochastic partial differential equation modulated by a stochastic volatility process. We analyze two classes of affine stochastic volatility models: (i) a Gaussian model governed by a finite-rank Wishart process, and (ii) a pure-jump affine model extending the Barndorff–-Nielsen–Shephard framework with state-dependent jumps in the covariance component. For both models, we derive conditions for the existence of exponential moments and develop semi-closed Fourier-based pricing formulas for vanilla call and put options written on forward price curves. Our approach allows for tractable pricing in models with infinitely many risk factors, thereby capturing maturity-specific and term structure risk essential in forward markets.
Keywords: option pricing, stochastic volatility models, affine processes, forward curve dynamics, commodity derivatives, Wishart processes, Fourier methods, HJM framework, exponential moments.
Asma Khedher is grateful for the financial support by the Research Foundation Flanders (FWO) under the grant FWO WOG W001021N
Jian He Sven Karbach Asma Khedher
[j.he2@uva.nl] [sven@karbach.org] [a.khedher@uva.nl]
1. Introduction
We consider a tractable framework for modeling forward price and forward rate curves in commodity and fixed-income markets, and develop semi-closed pricing formulas for the valuation of European vanilla options written on (potentially) the entire forward curve. Following the Heath–Jarrow–Morton–Musiela (HJMM) paradigm, originally proposed for interest-rate markets by [29] and later transferred to commodity markets by [11, 3], we describe forward prices directly in time-to-maturity coordinates. In this representation, the forward curve evolves according to a hyperbolic stochastic partial differential equation (SPDE) driven by a cylindrical Brownian motion whose covariance is itself modeled by a stochastic process that we call the instantaneous covariance process. We investigate two affine specifications of the covariance process: (i) An extension of the finite-rank Wishart process to Hilbert–Schmidt operators introduced in [14] and (ii) a state-dependent jump extension of the operator-valued Ornstein–Uhlenbeck process introduced in [16]. As both models are from the affine class, their associated Fourier-Laplace transforms admit semi-explicit expressions in terms of the solutions to a set of generalized Riccati equations, which allows for semi-closed pricing formulas for vanilla call and put options. The derivation of these pricing formulas, together with rigorous conditions for the existence of exponential moments and a numerical analysis, constitut the principal contribution of this article.
1.1. The HJMM framework
For any time $t\geq 0$ and maturity date $T\geq t$, the forward price $F(t,T)$ represents the agreed-upon price at time $t$ for a transaction occurring at the future date $T$. Utilizing the Musiela parametrization $x\equiv T-t$, forward prices become functions of time-to-maturity. We define the logarithmic forward prices as $f_{t}(x)\equiv\ln(F(t,t+x))$, where $x,t\geq 0$, and call the mapping $x\mapsto f_{t}(x)$ the (logarithmic) forward price curve at time $t\geq 0$.
The forward curve dynamics are described by the stochastic partial differential equation (SPDE):
$\begin{cases}\mathop{}\!\textnormal{d}f_{t}(x)&=\left(\frac{\partial}{\partial x}f_{t}(x)+g_{t}(x)\right)\mathop{}\!\textnormal{d}t+\sum_{i=1}^{d}\sigma_{t}^{(i)}(x)\mathop{}\!\textnormal{d}W_{t}^{(i)},\\ f_{0}(x)&=F(0,x),\end{cases}$ (1.1)
where $(W^{(i)}_{t})_{0\leq t\leq T}$, ${i=1},\ldots,{d}$, are independent Brownian motions. The drift term $g_{t}(x)$ satisfies a HJM no-arbitrage condition (see [9, Sec. 2.4] and Section 3 below), ensuring that each forward price process $(F(t,T))_{0\leq t\leq T}$ is a martingale under some risk-neutral measure $\mathbb{Q}$.
The functions $\{\sigma_{t}^{(i)}(x))_{0\leq t\leq T}$, ${i=1},\ldots,{d}$, are referred to as the instantaneous volatilities of the forward curve. They characterize the local covariance structure between forward prices of different maturities. The integer $d\in\mathbb{N}\cup\{+\infty\}$ is called the rank of the noise and corresponds to the number of independent risk factors driving the forward curves. If $d<\infty$, we refer to (1.1) as a finite-rank HJM model; otherwise, it is an infinite-rank HJM model.
1.2. Stochastic covariance models for forward curve dynamics
The diffusion term in (1.1) can be interpreted in a more abstract form as the stochastic integral of an operator-valued process $(\sigma_{t})_{0\leq t\leq T}$ acting on the increments of a cylindrical Brownian motion $(\mathrm{d}W_{t})_{0\leq t\leq T}$. We understand $(\sigma_{t})_{0\leq t\leq T}$ as the stochastic volatility operator, encoding the time-varying covariance structure of the forward curve dynamics. The functions $\sigma_{t}^{(i)}$ can therefore be interpreted as the principal components of $\sigma_{t}$, capturing the volatility structure along different directions in the underlying space of forward curves. The joint evolution of the forward curve and its stochastic volatility, represented by the process $(f_{t},\sigma_{t})_{0\leq t\leq T}$, thus defines an infinite-dimensional stochastic volatility model for forward curve dynamics.
The motivation for adopting a stochastic volatility framework stems from empirical evidence in commodity markets: volatility is time-varying, exhibits clustering, and is prone to abrupt spikes; see, e.g., [24, 6]. Classical models with constant volatility cannot capture these phenomena, whereas stochastic volatility models allow for a more realistic and flexible description of market behavior. Furthermore, to accurately capture volatility smiles in commodity option markets, incorporating stochastic volatility is essential. In contrast to approaches that perform a priori dimension reduction of the covariance, we do not restrict the rank of the volatility operator based on principal component analysis (PCA) or other variance-based criteria. Instead, we adopt a flexible, full-rank modeling approach, retaining all potential sources of risk. This is motivated by the observation that discarding components solely on the basis of explained variance can result in suboptimal choices for higher-order, non-linear functionals, such as those relevant for pricing, hedging, or risk management [12]. Note that even when the true rank is only moderately high, our methodology first emphasizes maximal expressiveness and generality. Finite-rank approximations are introduced a posteriori, and we study convergence of option prices as the rank increases. Empirical studies support this design choice: for instance, [1] and [34] demonstrate that power forward markets exhibit substantial idiosyncratic risk, and that high-dimensional noise is needed to capture the observed variability. In particular, the latter shows that even in PCA, around 10 factors are typically required to explain the majority of the variance, underlining the importance of accommodating high-dimensional dynamics from the outset. Here our function-valued formulation helps, as it is essentially non-parametric, relying on the full initial forward curve and the full covariance operator as model inputs. The covariance structure is encoded via a positive kernel operator, which can capture empirical features with a minimal set of hyperparameters and enables direct calibration to market data, including option prices and historical term structure dynamics.
The stochastic volatility models considered in this paper belong to the class of affine processes, which play a central role in modern financial modeling due to their analytical tractability. Affine processes satisfy the so-called affine transform formula, which expresses the cumulant generating function as an affine function of the initial state and the solution of a system of generalized Riccati ordinary differential equations (ODEs). A key advantage of this framework is the quasi-explicit availability of the Fourier–Laplace transform, which enables efficient and accurate pricing of options and other derivatives via Fourier inversion methods; see [10, 22, 23]. The extension of affine modeling techniques to infinite-dimensional settings has been the subject of extensive research; see, e.g., [37, 19, 5, 8, 30, 31, 28, 17, 2]. In parallel, robust statistical methods have been developed for the estimation and analysis of stochastic volatility and covariation in term structure models. These approaches address both the functional nature of the data and the dynamic consistency required for sound modeling; see [39, 38].
1.3. Contribution
In this paper, we contribute to this growing literature by developing a general pricing framework for European options written on forward contracts under function-valued affine stochastic volatility. We formulate the forward curve dynamics within the HJMM framework and model the instantaneous covariance as an operator-valued affine process. Two concrete model classes are considered: a Gaussian specification driven by an infinite-dimensional Wishart process, and a pure-jump model based on an extension of an operator-valued Ornstein–Uhlenbeck process to accommodate state-dependent jumps. The inclusion of jumps in the volatility process enhances the model’s ability to reflect sudden market shocks, which commonly arise from supply disruptions, geopolitical tensions, or abrupt changes in demand; see, e.g. [35]. In both settings, we rigorously extend the affine transform formula to a complex strip in an infinite-dimensional Hilbert space and establish conditions under which exponential moments exist in real and complex domains. In particular, we characterize the maximal domain of the moment generating function using generalized Riccati equations in Banach spaces, building on the analytic machinery introduced in [36], and extending prior finite-dimensional results such as those in [33].
Based on these results, we derive semi-closed-form pricing formulas for vanilla call and put options, which enable efficient numerical evaluation via Fourier inversion. We validate the accuracy and robustness of this pricing methodology through a comparison with Monte Carlo simulation. The numerical results show that both the moment-generating function (MGF) and option prices derived via the affine transform are consistent with those obtained from Monte Carlo simulations, see Figure 1, 2 and 3 below. This comparison not only validates the accuracy of our approach but also illustrates its computational advantages. For example, in the case of the infinite-dimensional Wishart-type volatility model, direct simulation is challenging due to the absence of a closed-form expression for the square root of the operator-valued volatility process. Approximating this square root requires advanced numerical techniques, making Monte Carlo simulation computationally intensive. Similarly, in the case of the infinite-dimensional pure-jump affine volatility model, Monte Carlo methods demand significant computational resources, both in memory and runtime, due to the complexity of the jump structure. By contrast, the affine structure allows for efficient computation of key quantities such as the MGF and derivative prices without resorting to discretization of the infinite-dimensional dynamics. This leads to considerable gains in computational speed and accuracy. These advantages make affine stochastic volatility models particularly suitable for practical applications. In particular, they facilitate the calibration to implied volatility surfaces and option price curves observed in fixed-income or commodity markets, which is a notoriously delicate task. Our results demonstrate that, despite the infinite-dimensional nature of the models, their tractability and analytical structure make them powerful tools for capturing the rich dynamics in forward and futures markets.
1.4. Layout of the article
The remainder of the paper is organized as follows. In Section 2, we introduce the modeling framework for forward curve dynamics using the HJMM approach and define the class of infinite-dimensional stochastic volatility models under consideration. Sections 2.1 and 2.2 present two concrete affine specifications for the volatility process: a pure-jump model based on an extension of an operator-valued Ornstein–Uhlenbeck process, and a Gaussian model driven by an infinite-dimensional Wishart process. In Section 3, we provide semi-closed-form pricing formulas for European call and put options on forwards, obtained via Fourier inversion. Section 4 presents a numerical study comparing the affine pricing formulas with Monte Carlo simulation methods. We highlight the accuracy, robustness, and computational efficiency of the affine approach, especially in settings where simulation becomes numerically challenging. Finally, in Section 5, we study the existence of the exponential moments.
1.5. Notation
For $(X,\tau)$ a topological space and $S\subset X$ we let $\mathcal{B}(S)$ denote the Borel-$\sigma$-algebra generated by the relative topology on $S$. We denote by $C^{k}([0,T];S)$ the space of $S$-valued $k$-times continuously differentiable functions on $[0,T]$. Throughout this article we fix a separable, infinite-dimensional real Hilbert space $(H,\langle\cdot,\cdot\rangle_{H})$. The complexification of a real Hilbert space is the complex Hilbert space $H^{\mathbb{C}}=H\oplus_{\mathbb{R}}\mathrm{i}H=\{x+\mathrm{i}y\colon x,y\in H\}$ equipped with the complex inner product $\langle x+\mathrm{i}y,u+\mathrm{i}v\rangle_{H^{\mathbb{C}}}=\langle x,u\rangle_{H}-\langle y,v\rangle_{H}+\mathrm{i}\langle y,u\rangle_{H}+\mathrm{i}\langle x,v\rangle_{H}$, for all $x,y,u,v\in H$. For $z=x+\mathrm{i}y\in H^{\mathbb{C}}$, we say that $x=\Re(z)$ and $y=\mathcal{I}(z)$ are the real and imaginary parts of $z$. Moreover, $\bar{z}=\Re(z)-\mathrm{i}\,\mathcal{I}(z)$ denotes the conjugate of $z$.
The space of bounded linear operators from $H$ to $H$ is denoted by $\mathcal{L}(H)$. The space of compact operators from $H$ to $H$ is denoted by $(K(H),\left\|\cdot\right\|_{\mathcal{L}(H)})$. The adjoint of an operator $A\in\mathcal{L}(H)$ is denoted by $A^{*}$, i.e., $\langle Ah,g\rangle_{H}=\langle h,A^{*}g\rangle_{H}$ for all $h\in H$. For $A\in\mathcal{L}(H^{\mathbb{C}})$ we define the transpose $A^{T}\in\mathcal{L}(H^{\mathbb{C}})$ by $\langle A^{T}h,g\rangle_{H^{\mathbb{C}}}=\langle A\bar{g},\bar{h}\rangle_{H^{\mathbb{C}}}$, $h,g\in H^{\mathbb{C}}$. Note that in contrast to $A^{*}$, the operator $A^{T}$ corresponds to the transpose without conjugation. Note that any $A\in\mathcal{L}(H)$ extends in a trivial and norm-conserving way to an operator $\tilde{A}\in\mathcal{L}(H^{\mathbb{C}})$ by setting $\tilde{A}(h+ig)=Ah+iAg$, $h,g\in H$.
For all $p\in[1,\infty)$ let $(\mathcal{L}_{p}(H),\left\|\cdot\right\|_{\mathcal{L}_{p}(H)})$ be the Banach space of Schatten class operators from $H$ to $H$, i.e.,
$\mathcal{L}_{p}(H)=\Big{\{}A\in K(H)\colon\sum_{\lambda\in\sigma(A^{*}A)}\lambda^{p/2}<\infty\Big{\}},$ (1.2)
and $\|A\|_{\mathcal{L}_{p}(H)}^{p}=\sum_{\lambda\in\sigma(A^{*}A)}\lambda^{p/2}$. In particular, we let $\mathcal{L}_{1}(H)$ and $\mathcal{L}_{2}(H)$ denote respectively the space of trace class operators and the space of Hilbert-Schmidt operators on $H$. Recall that $\mathcal{L}_{2}(H)$ is a Hilbert space when endowed with the inner product
$\langle A,B\rangle_{\mathcal{L}_{2}(H)}=\sum_{n=1}^{\infty}\langle Ae_{n},Be_{n}\rangle_{H}.$
For notational brevity, we reserve $\langle\cdot,\cdot\rangle$ to denote the inner product on $\mathcal{L}_{2}(H)$, and $\|\cdot\|_{2}$ for the norm induced by $\langle\cdot,\cdot\rangle$. Moreover, we reserve $\|\cdot\|_{1}$ for the norm $\|\cdot\|_{\mathcal{L}_{1}(H)}$. We recall that the trace of $A\in\mathcal{L}_{1}(H)$ is defined by
$\operatorname{Tr}(A)=\sum_{n\in\mathbb{N}}\langle Ah_{n},h_{n}\rangle_{H}\in\mathbb{C},$ (1.3)
where $(h_{n})_{n\in\mathbb{N}}$ is an orthonormal basis for $H$; $\operatorname{Tr}(A)$ does not depend on the choice of the orthonormal basis.
Writing $V^{\prime}$ for the dual of a Banach space $V$, we recall (see, e.g., [13, Section 19]) that the dual space of trace class operators satisfies $(\mathcal{L}_{1}(H_{1}))^{\prime}\simeq\mathcal{L}(H)$ under the duality paring
$\langle A,B\rangle_{\mathcal{L}(H),\mathcal{L}_{1}(H)}=\operatorname{Tr}(B^{*}A),\quad A\in\mathcal{L}(H),\,B\in\mathcal{L}_{1}(H).$ (1.4)
We let $\mathcal{S}(H)$, $\mathcal{S}_{1}(H)$ denote the (closed) subspaces of $\mathcal{L}(H)$, $\mathcal{L}_{1}(H)$ consisting of all operators that are self-adjoint and we let $\mathcal{S}^{+}(H)$ and $\mathcal{S}_{1}^{+}(H)$ denote the (closed) subsets of $\mathcal{S}(H)$ and $\mathcal{S}_{1}(H)$ consisting of all self-adjoint operators $A$ satisfying $\sigma(A)\subset[0,\infty)$. Moreover, if $A\in\mathcal{S}(H)$, then the extended operator $\tilde{A}\in\mathcal{L}(H^{\mathbb{C}})$ fulfills $\tilde{A}^{T}=\tilde{A}=\tilde{A}^{*}$, in particular, $\tilde{A}\in\mathcal{S}(H^{\mathbb{C}})$ (from now on we do not distinguish between $A$ and $\tilde{A}$). In this article we will encounter the set $\mathcal{S}^{+}(H)\oplus\mathrm{i}\mathcal{S}(H)\subseteq\mathcal{L}(H^{\mathbb{C}})$, which denotes the operators $A\in\mathcal{L}(H^{\mathbb{C}})$ for which there exist (necessarily unique) $A_{1}\in\mathcal{S}^{+}(H)$, $A_{2}\in\mathcal{S}(H)$ such that $A=A_{1}+\mathrm{i}A_{2}$. The set $\mathcal{S}(H)\oplus\mathrm{i}\mathcal{S}(H)$ is defined analogously. Note that if $A\in\mathcal{S}(H)\oplus\mathrm{i}\mathcal{S}(H)$, then $A^{T}=A$.
A nonempty subset $K$ of a vector space is called a wedge if $K+K\subseteq K$ and $\alpha K\subseteq K$ for all $\alpha\geq 0$, if moreover $K\cap(-K)=\{0\}$ then we call $K$ a cone. A cone $K$ in a vector space $X$ induces a partial ordering: we write $x\leq_{K}y$ if $y-x\in K$ (and $x\geq_{K}y$ if $x-y\in K$). For a Banach space $V$, if $K\subset V$ is a wedge, we define the dual of $K$ by
[…]
and we say that $K$ is self-dual if $K=K^{*}$. We define $\mathcal{H}$ to be the space of all self-adjoint Hilbert-Schmidt operators on $H$ and $\mathcal{H}^{+}$ to be the cone of all positive operators in $\mathcal{H}$:
$\mathcal{H}:=\{A\in\mathcal{L}_{2}(H)\colon A=A^{*}\},\ \text{and}\ \mathcal{H}^{+}:=\{A\in\mathcal{H}\colon\langle Ah,h\rangle_{H}\geq 0\text{ for all }h\in H\}.$
Note that $\mathcal{H}$ is a closed subspace of $\mathcal{L}_{2}(H)$, and that $\mathcal{H}^{+}$ is a self-dual cone in $\mathcal{H}$. For $a,b\in H$, we let $a\otimes b$ be the linear operator defined by $a\otimes b(h)=\langle a,h\rangle_{H}b$ for every $h\in H$. Note that $a\otimes a\in\mathcal{H}^{+}$ for every $a\in H$ and when space is scarce, we shall write $a^{\otimes 2}\coloneqq a\otimes a$. Finally, throughout this article we let $\chi\colon\mathcal{H}\rightarrow\mathcal{H}$ denote the truncation function given by $\chi(x)=x\mathbbm{1}_{\{\|x\|_{2}\leq 1\}}$.
2. Infinite-Dimensional Affine Stochastic Volatility Models
In this section, we present two affine stochastic volatility models in an infinite-dimensional function-valued setting. In both cases, the underlying forward-curve process solves the HJMM SPDE (1.1) in some separable Hilbert space $H$ of forward curves. The key distinction between the two models lies in the specification of their instantaneous covariance processes $(\sigma_{t})_{t\geq 0}$.
2.1. The affine pure-jump stochastic volatility model
We begin by formally introducing the two instantaneous covariance processes and their respective joint Fourier-Laplace transform.
2.1.1. A pure-jump $\mathcal{H}^{+}$-valued instantaneous covariance process
To introduce our pure-jump covariance process, we need the following admissibility conditions that we impose on the parameter of the process:
Assumption $\mathfrak{A}$.
An admissible parameter set $(b,B,m,\mu)$ consists of
  1. (i)
    a measure $m\colon\mathcal{B}(\mathcal{H}^{+}\setminus\{0\})\to[0,\infty]$ such that
    1. (a)
      $\int_{\mathcal{H}^{+}\setminus\{0\}}\|\xi\|_{2}^{2}\,m(\mathop{}\!\textnormal{d}\xi)<\infty$ and
    2. (b)
      $\int_{\mathcal{H}^{+}\setminus\{0\}}|\langle\chi(\xi),h\rangle|\,m(\mathop{}\!\textnormal{d}\xi)<\infty$ for all $h\in\mathcal{H}$ and there exists an element $I_{m}\in\mathcal{H}$ such that $\langle I_{m},h\rangle=\int_{\mathcal{H}^{+}\setminus\{0\}}\langle\chi(\xi),h\rangle\,m(\mathop{}\!\textnormal{d}\xi)$ for every $h\in\mathcal{H}$ ;
  2. (ii)
    a vector $b\in\mathcal{H}$ such that
    $\langle b,v\rangle-\int_{\mathcal{H}^{+}\setminus\{0\}}\langle\chi(\xi),v\rangle\,m(\mathop{}\!\textnormal{d}\xi)\geq 0\,\quad\text{for all}\;v\in\mathcal{H}^{+}\,;$ (2.1)
  3. (iii)
    a $\mathcal{H}^{+}$-valued measure $\mu\colon\mathcal{B}(\mathcal{H}^{+}\setminus\{0\})\rightarrow\mathcal{H}^{+}$ such that
    $\int_{\mathcal{H}^{+}\setminus\{0\}}\langle\chi(\xi),u\rangle\frac{\langle\mu(\mathop{}\!\textnormal{d}\xi),x\rangle}{\|\xi\|_{2}^{2}}<\infty,$
    for all $u,x\in\mathcal{H}^{+}$ satisfying $\langle u,x\rangle=0$ ;
  4. (iv)
    an operator $B\in\mathcal{L}(\mathcal{H})$ with adjoint $B^{*}$ satisfying
    $\left\langle B^{*}(u),x\right\rangle-\int_{\mathcal{H}^{+}\setminus\{0\}}\langle\chi(\xi),u\rangle\frac{\langle\mu(\mathop{}\!\textnormal{d}\xi),x\rangle}{\|\xi\|_{2}^{2}}\geq 0,$
    for all $x,u\in\mathcal{H}^{+}$ satisfying $\langle u,x\rangle=0$.
In this paper, we consider affine pure-jump covariance processes that are either driven by a pure jump Lévy process (nullifying the state-dependent jump parts), or assume that the jump-process is of finite-variation, which is due to the existence of exponential moments in these two settings (see Section 5 below).
Assumption $\mathfrak{B}$.
Assume that $(b,B,m,\mu)$ is an admissible parameter such that $\mu(\mathop{}\!\textnormal{d}\xi)$ takes values in $\mathcal{L}_{1}(H)\cap\mathcal{H}^{+}$. Moreover, assume that either one of the following two cases holds:
  1. (i)
    $\mu=0$,
  2. (ii)
    $\int_{\mathcal{H}^{+}\cap\{\|\xi\|_{2}\leq 1\}}\|\xi\|_{2}\,m(\mathop{}\!\textnormal{d}\xi)<\infty\,$ and $\int_{\mathcal{H}^{+}\cap\{\|\xi\|_{2}\leq 1\}}\|\xi\|_{2}^{-1}\langle y,\mu(\mathop{}\!\textnormal{d}\xi)\rangle<\infty$, $\forall y\in\mathcal{H}^{+}.$
Given an admissible parameter set $(b,B,m,\mu)$ that satisfies Assumptions Assumption $\mathfrak{A}$ and Assumption $\mathfrak{B}$, we conclude from [15, Theorem 2.8] that there exists a probabilistically and analytically weak solution to
$\mathop{}\!\textnormal{d}\mathsf{Y}_{t}$ $=\Big{(}b+B(\mathsf{Y}_{t})+\int_{\mathcal{H}^{+}\cap\{\|\xi\|_{2}>1\}}\xi\,M(\mathsf{Y}_{t},\mathop{}\!\textnormal{d}\xi)\Big{)}\mathop{}\!\textnormal{d}t+\mathop{}\!\textnormal{d}J_{t},\qquad t\geq 0\,$, (2.2)
$\mathsf{Y}_{0}$ $=\mathsf{y}\in\mathcal{H}^{+}\,$,
where $(J_{t})_{t\geq 0}$ is a purely-discontinuous square integrable $\mathcal{H}$-valued martingale with jump intensity measure depending on the state $\mathsf{Y}_{t}$ in affine manner as
$M(y,\mathop{}\!\textnormal{d}\xi):=m(\mathop{}\!\textnormal{d}\xi)+\frac{\langle\mu(\mathop{}\!\textnormal{d}\xi),y\rangle}{\|\xi\|_{2}^{2}}.$ (2.3)
Notice that $(\mathsf{Y}_{t})_{t\geq 0}$ is a square integrable $\mathcal{H}^{+}$-valued Markovian semimartingale and (2.2) amounts to the semimartingale representation of $(\mathsf{Y}_{t})_{t\geq 0}$, see [16, Proposition 2.6] for the details.
2.1.2. The affine pure-jump joint stochastic volatility model
The first stochastic volatility model we consider takes as instantaneous volatility process the square root of the pure-jump $\mathcal{H}^{+}$-valued process from (2.2). We denote the forward price dynamics by $(X_{t})_{t\geq 0}$, governed by the HJMM–SPDE (1.1), but formulated in a Hilbert space $H$ as follows. Fix an initial value $(x,\mathsf{y})$ with $\mathsf{y}\in\mathcal{H}^{+}$ and $x\in H$, and model parameters $(b,B,m,\mu,D,\Upsilon,\mathcal{A})$. We consider the coupled SDE
$\begin{cases}\mathop{}\!\textnormal{d}X_{t}=\big{(}\mathcal{A}X_{t}+D^{1/2}\mathsf{Y}_{t}D^{1/2}\Upsilon\big{)}\,\mathop{}\!\textnormal{d}t+D^{1/2}\mathsf{Y}_{t}^{1/2}\,\mathop{}\!\textnormal{d}W_{t},\\ \mathop{}\!\textnormal{d}\mathsf{Y}_{t}=\Big{(}b+B(\mathsf{Y}_{t})+\int_{\mathcal{H}^{+}\cap\{\|\xi\|_{2}>1\}}\xi\,M(\mathsf{Y}_{t},\mathop{}\!\textnormal{d}\xi)\Big{)}\mathop{}\!\textnormal{d}t+\mathop{}\!\textnormal{d}J_{t},\quad t\geq 0,\end{cases}$ (2.4)
where $(W_{t})_{t\geq 0}$ and $(J_{t})_{t\geq 0}$ are independent. Here $(b,B,m,\mu)$ is an admissible parameter set in the sense of Assumption Assumption $\mathfrak{A}$, $(\mathcal{A},\operatorname{dom}(\mathcal{A}))$ denotes the first-derivative (in the maturity variable) operator on the Hilbert space $H$, $D\in\mathcal{L}_{1}(H)$, and $\Upsilon\in H$; see Section 3 for details.
The existence of a (stochastically) weak solution to (2.4) in $H\times\mathcal{H}$ follows from [16, Proposition 2.6 and Lemma 2.8]. As in [16], and motivated by applications, we include the linear term $D^{1/2}\mathsf{Y}_{t}D^{1/2}\Upsilon$ in the drift of $(X_{t})_{t\geq 0}$. We refer to (2.4) as the pure-jump affine stochastic volatility model.
2.1.3. Riccati equations associated with the pure-jump affine volatility model
Next we aim to derive the Laplace transform of our joint stochastic volatility model. The Laplace will be computed in terms of a solution to ordinary differential equations known as Riccati equations. We start by introducing the Riccati equations associated with our introduced model.
Let $(b,B,m,\mu)$ be an admissible parameter set satisfying Assumptions Assumption $\mathfrak{A}$. Recall the measure $M$ in (2.3) and define
$\mathcal{O}=\left\{u\in\mathcal{H}\colon\int_{\mathcal{H}^{+}\cap\left\{\|\xi\|_{2}>1\right\}}\mathop{}\!\textnormal{e}^{-\langle\xi,u\rangle}\,M(y,\mathop{}\!\textnormal{d}\xi)<\infty,\quad\forall y\in\mathcal{H}^{+}\right\}.$
We note that $\mathcal{O}$ is a convex subset in $\mathcal{H}$ and since $\mathcal{H}^{+}\subseteq\mathcal{O}$, it is non-empty. Moreover, we assume that $M(x,\mathop{}\!\textnormal{d}\xi)$ is such that $\mathcal{O}$ is an open subset of $\mathcal{H}$. Define ${\sf F}\colon\mathcal{O}\to\mathbb{R}$ and ${\sf R}\colon H\times\mathcal{O}\to\mathcal{H}$, respectively as
${\sf F}(u)$ $=\langle\tilde{b},u\rangle-\int_{\mathcal{H}^{+}\setminus\{0\}}\big{(}\mathop{}\!\textnormal{e}^{-\langle\xi,u\rangle}-1\big{)}m(\mathop{}\!\textnormal{d}\xi),$ (2.5)
${\sf R}(h,u)$ $=\tilde{B}^{*}(u)-\tfrac{1}{2}(D^{1/2}h)^{\otimes 2}-D^{1/2}h\otimes D^{1/2}\Upsilon-\int_{\mathcal{H}^{+}\setminus\{0\}}\big{(}\mathop{}\!\textnormal{e}^{-\langle\xi,u\rangle}-1)\frac{\mu(\mathop{}\!\textnormal{d}\xi)}{\|\xi\|_{2}^{2}},$ (2.6)
where $\tilde{b}\coloneqq b-\int_{\mathcal{H}^{+}\setminus\{0\}}\chi(\xi)\,m(\mathop{}\!\textnormal{d}\xi)$ and for all $v\in\mathcal{H}^{+}$
$\tilde{B}(v)\coloneqq B(v)-\int_{\mathcal{H}^{+}\setminus\{0\}}\chi(\xi)\frac{\langle v,\mu(\mathop{}\!\textnormal{d}\xi)\rangle}{\|\xi\|_{2}^{2}}.$
Note that $\tilde{b}$ and $\tilde{B}$ are well defined due to the finite-variation Assumption Assumption $\mathfrak{B}$.
Definition 2.1.
Let $T\geq 0$ and $u=(u_{1},u_{2})\in H\times\mathcal{O}$. Consider the following extended generalized Riccati equations:
$\,\frac{\partial\sf\sf P}{\partial t}(t,u)$ $={\sf F}({\sf q}_{2}(t,u)),$  $\,0<t\leq T,\quad{\sf P}(0,u)=0,$ (2.7a)
$\,{\sf q}_{1}(t,u)$ $=u_{1}+\mathcal{A}^{*}\left(\int_{0}^{t}{\sf q}_{1}(s,u)\mathop{}\!\textnormal{d}s\right),$  $\,0<t\leq T,\quad{\sf q}_{1}(0,u)=u_{1},$ (2.7b)
$\,\frac{\partial{\sf q}_{2}}{\partial t}(t,u)$ $={\sf R}({\sf q}_{1}(t,u),{\sf q}_{2}(t,u)),$  $\,0<t\leq T,\quad{\sf q}_{2}(0,u)=u_{2}.$ (2.7c)
For $u=(u_{1},u_{2})\in H\times\mathcal{O}$ and $T\geq 0$, we say that $({\sf P}(\cdot,u),{\sf q}_{1}(\cdot,u),{\sf q}_{2}(\cdot,u))\colon[0,T]\rightarrow\mathbb{R}\times H\times\mathcal{H}$ is a mild solution to (2.7a)-(2.7c) if ${\sf P}(\cdot,u)\in C^{1}([0,T],\mathbb{R})$, ${\sf q}_{1}(\cdot,u)\in C([0,T],H)$ and ${\sf q}_{2}(\cdot,u)\in C^{1}([0,T],\mathcal{O})$ satisfies (2.7a)-(2.7c).
Let $\mathcal{H}^{\mathbb{C}}$ denote the complexification of $\mathcal{H}$. We introduce the following notation
$S(\mathcal{O})\coloneqq\left\{u\in\mathcal{H}^{\mathbb{C}}\colon\,\Re(u)\in\mathcal{O}\right\}.$
Note that for every $u\in S(\mathcal{O})$ we have
$\int_{\mathcal{H}^{+}\cap\left\{\|\xi\|_{2}>1\right\}}|\mathop{}\!\textnormal{e}^{-\langle\xi,u\rangle}|M(y,\mathop{}\!\textnormal{d}\xi)=\int_{\mathcal{H}^{+}\cap\left\{\|\xi\|_{2}>1\right\}}\mathop{}\!\textnormal{e}^{-\langle\xi,\Re(u)\rangle}M(y,\mathop{}\!\textnormal{d}\xi)<\infty\,,\quad\forall y\in\mathcal{H}^{+}\,.$
Definition 2.2.
We shall consider $\mathsf{F}$ and $\mathsf{R}$ as functions from $S(\mathcal{O})$ to $\mathbb{C}$ resp. from $H^{\mathbb{C}}\times S(\mathcal{O})$ to $\mathcal{H}^{\mathbb{C}}$. We then consider the following complex generalized Riccati equations:
$\,\frac{\partial{\mathsf{\phi}}}{\partial t}(t,u)$ $=\mathsf{F}(\psi_{2}(t,u)),\quad 0<t\leq T,\quad\phi(0,u)=0,$ (2.8a)
$\,\psi_{1}(t,u)$ $=u_{1}+\mathcal{A}^{*}\int_{0}^{t}\psi_{1}(s,u)\mathop{}\!\textnormal{d}s,\quad 0<t\leq T,\quad\psi_{1}(0,u)=u_{1}\,$, (2.8b)
$\,\frac{\partial\psi_{2}}{\partial t}(t,u)$ $=\mathsf{R}(\psi_{1}(t,u),\psi_{2}(t,u)),\quad 0<t\leq T,\quad\psi_{2}(0,u)=0.$ (2.8c)
Let $T\geq 0$, $u=(u_{1},0)$, for $u_{1}\in H^{\mathbb{C}}$. Analogously to the extended generalized Riccati equations (2.7a)-(2.7c), we say that the map $(\phi(\cdot,u),\psi_{1}(\cdot,u),\psi_{2}(\cdot,u))$ is a solution to equations (2.8a)-(2.8c) whenever $\phi(\cdot,u)\in C^{1}([0,T],\mathbb{C}),\psi_{1}(\cdot,u)\in C([0,T],H^{\mathbb{C}}),\psi_{2}(\cdot,u)\in C^{1}([0,T],S(\mathcal{O}))$ satisfy (2.8a)-(2.8c).
2.1.4. Fourier-Laplace transforms of the pure-jump affine volatility model
From Proposition 5.5 and following similar derivations as in [33, Theorem 2.26 and Section 5.3], we deduce the following result:
Theorem 2.3.
Let $(b,B,m,\mu)$ be an admissible parameter set satisfying Assumptions Assumption $\mathfrak{A}$ and Assumption $\mathfrak{B}$, let $(X,\sf Y)$ be the stochastic volatility model (2.4). Let $u=(u_{1},u_{2})\in H^{\mathbb{C}}\times\mathcal{S}(\mathcal{O})$. Let $(\sf P,\sf q_{1},\sf q_{2})$ be the mild solution to the extended Riccati equations (2.7a)-(2.7c) on $[0,T]$ with initial value $(\Re(u_{1}),\Re(u_{2}))$, the existence of which is guaranteed by Proposition 5.2. Then it holds $\mathbb{E}\left[|\mathop{}\!\textnormal{e}^{\langle X_{t},u_{1}\rangle_{H^{\mathbb{C}}}}|\right]<\infty$, and
$\mathbb{E}\left[\mathop{}\!\textnormal{e}^{\langle X_{T},u_{1}\rangle_{H^{\mathbb{C}}}}\mid\mathcal{F}_{t}\right]=\mathop{}\!\textnormal{e}^{-\phi(t,u_{1},0)+\langle X_{t},\psi_{1}(T-t,u_{1},0)\rangle_{H^{\mathbb{C}}}-\langle Y_{t},\psi_{2}(T-t,u_{1},0)\rangle_{\mathcal{H}^{\mathbb{C}}}},$
where $(\phi(\cdot,u_{1},0),\psi_{1}(\cdot,u_{1},0),\psi_{2}(\cdot,u_{1},0))$ is the unique solution to (2.8a)-(2.8c) on $[0,T]$ the existence of which is guaranteed by Proposition 5.5.
2.2. The infinite-dimensional Wishart stochastic volatility model
As in the case of the pure-jump stochastic volatility model, we start by introducing the instantaneous covariance process that defines a Gaussian infinite-dimensional stochastic volatility model.
2.2.1. Infinite-dimensional Wishart process
Let $\mathbb{A}\in\mathcal{L}(H)$, Let $Q\in\mathcal{S}^{+}(H)$, let $n\in\mathbb{N}$, and let $y\in\mathcal{S}_{1}^{+}(H)$ be of rank at most $n$. Moreover, assume that
$\int_{0}^{t}\|e^{s\mathbb{A}}\sqrt{Q}\|^{2}_{2}\,\mathop{}\!\textnormal{d}s<\infty\,$,
for all $t\geq 0$. Consider the following stochastic differential equation
$\mathop{}\!\textnormal{d}\mathcal{Y}_{t}$ $=nQ\,\mathop{}\!\textnormal{d}t+\mathcal{Y}_{t}\mathbb{A}\mathop{}\!\textnormal{d}t+\mathbb{A}^{*}\mathcal{Y}_{t}\,\mathop{}\!\textnormal{d}t+\sqrt{\mathcal{Y}}_{t}\,\mathop{}\!\textnormal{d}B_{t}\sqrt{Q}+\sqrt{Q}\mathop{}\!\textnormal{d}B_{t}^{*}\sqrt{\mathcal{Y}}_{t}\,,\quad t\geq 0\,$,
$\mathcal{Y}_{0}$ $=y\,$, (2.9)
where $(B_{t})_{t\geq 0}$ is an $\mathcal{L}_{2}(H)$-cylindrical Brownian motion. As shown in [14, Theorem 2.1], equation (2.2.1) admits a probabilistically and analytically weak solution. In fact, the existence result in [14, Theorem 2.1] holds more generally when $\mathbb{A}$ is an unbounded operator generating a strongly-continuous semigroup.
Furthermore, [14, Proposition 4.7 and Corollary 4.11] provide a characterization of finite-rank $\mathcal{S}_{1}^{+}(H)$-valued Wishart processes. In this paper, however, we restrict our attention to the case of a bounded operator $\mathbb{A}$, since under this assumption equation (2.2.1) can naturally serve as a model for the instantaneous covariance process as we argue next.
2.2.2. The infinite-dimensional Wishart stochastic volatility model
Let $n\in\mathbb{N}$, let $y\in\mathcal{S}_{1}^{+}(H)$ be of rank at most $n$, and let $(W_{t})_{t\geq 0}$ be an $H$-cylindrical Brownian motion. Let $x,\Upsilon\in H$ and $D\in\mathcal{S}_{1}^{+}(H)$, and let $(X_{t})_{t\geq 0}$ be given as the solution to the HJMM–SPDE (1.1).
In this second model, the instantaneous covariance process is given by the square root of the Wishart process (2.2.1). For parameters $(\mathbb{A},Q,D,\Upsilon,\mathcal{A})$ and initial value $(x,y)$, we consider the coupled SDE
$\begin{cases}\mathop{}\!\textnormal{d}X_{t}=\big{(}\mathcal{A}X_{t}+D^{1/2}\mathcal{Y}_{t}D^{1/2}\Upsilon\big{)}\,\mathop{}\!\textnormal{d}t+D^{1/2}\mathcal{Y}_{t}^{1/2}\,\mathop{}\!\textnormal{d}W_{t},\\[4.0pt] \mathop{}\!\textnormal{d}\mathcal{Y}_{t}=\big{(}nQ+\mathcal{Y}_{t}\mathbb{A}+\mathbb{A}^{*}\mathcal{Y}_{t}\big{)}\,\mathop{}\!\textnormal{d}t+\sqrt{\mathcal{Y}_{t}}\,\mathop{}\!\textnormal{d}B_{t}\sqrt{Q}+\sqrt{Q}\,\mathop{}\!\textnormal{d}B_{t}^{*}\sqrt{\mathcal{Y}_{t}},\quad t\geq 0,\end{cases}$ (2.10)
where $(W_{t})_{t\geq 0}$ and $(B_{t})_{t\geq 0}$ are independent. The existence of a (stochastically) weak solution to (2.10) in $H\times\mathcal{S}_{1}^{+}(H)$ follows from [16, Lemma 2.8] and [14, Theorem 2.1].
Following [16], and motivated by applications, we include the linear term $D^{1/2}\mathcal{Y}_{t}D^{1/2}\Upsilon$ in the drift of $(X_{t})_{t\geq 0}$. This addition does not affect well-posedness or the affine property; see, for example, the proof of [16, Lemma 2.8]. We refer to (2.10) as the infinite-dimensional Wishart stochastic volatility model.
2.2.3. Riccati equations associated with the infinite-dimensional Wishart stochastic volatility model
Let $n\in\mathbb{N}$, $Q,D\in\mathcal{S}_{1}^{+}(H)$, $\Upsilon\in H$, and $\mathbb{A}\in\mathcal{L}(H)$. Define $F\colon\mathcal{S}_{1}(H)\to\mathbb{R}$ and $R\colon H\times\mathcal{S}_{1}(H)\to\mathcal{L}_{1}(H)$ by
$F(u)$ $=n\,\operatorname{Tr}(Qu),$ (2.11)
$R(h,u)$ $=u\mathbb{A}^{*}+\mathbb{A}u-\tfrac{1}{2}\big{(}D^{1/2}h\big{)}^{\otimes 2}-D^{1/2}h\otimes D^{1/2}\Upsilon$
$\quad-\tfrac{1}{2}\big{(}u+u^{T}\big{)}Q\big{(}u+u^{T}\big{)}.$ (2.12)
As in the affine pure-jump case, the functions $F$ and $R$ give rise to a system of operator-valued Riccati-type equations. Since the operators $(\mathcal{Y}_{t})_{t\geq 0}$ of the Wishart process are, by construction, of finite rank, we adopt a numerically convenient simplification: we approximate the model by a matrix-valued Wishart process whose rank remains constant over time. This reduces the infinite-dimensional problem (at least in the covariance dimensions) to a more tractable finite-dimensional one while preserving the essential structure of the model.
Let $(H_{n})_{n\in\mathbb{N}}$ denote an increasing sequence of finite-dimensional subspaces such that $\overline{\bigcup_{n}H_{n}}=H$. Let $\{e_{1},\ldots,e_{n}\}$ be an orthonormal basis of $H_{n}$ and define the orthogonal projection $\Pi_{n}:H\to H_{n}$. We denote by $\mathcal{L}_{n}:=\mathcal{L}(H_{n})\cong\mathbb{R}^{n\times n}$ the space of bounded linear operators on $H_{n}$ and by $\mathcal{S}_{n}$ the space of real symmetric $n\times n$ matrices. Define the finite-rank counterparts $F_{n}\colon\mathcal{S}_{n}\to\mathbb{R}$ and $R_{n}\colon H\times\mathcal{S}_{n}\to\mathcal{L}(H_{n})$ of the nonlinear maps $F$ and $R$ introduced in (2.11) and (2.12) via
$F_{n}(X):=F(\iota_{n}X\iota_{n}^{*}),\qquad R_{n}(y,X):=\Pi_{n}R(y,\iota_{n}X\iota_{n}^{*})\Pi_{n}^{*},$ (2.13)
where $\iota_{n}:\mathbb{R}^{n}\hookrightarrow H$ is the canonical injection.
Definition 2.4.
Let $F_{n}$ and $R_{n}$ be as in (2.13). Then, for $u=(u_{1},u_{2})\in H\times\mathcal{S}_{n}$ and $T\geq 0$, the finite-rank Riccati system is given by the system of ordinary differential equations
$\frac{\partial P_{n}}{\partial t}(t,u)$ $=F_{n}(q_{2,n}(t,u)),\quad$ $P_{n}(0,u)=0,$ (2.14a)
$q_{1}(t,u)$ $=u_{1}+\mathcal{A}^{*}\int_{0}^{t}q_{1}(s,u)\,\mathrm{d}s,\quad$ $q_{1}(0,u)=u_{1},$ (2.14b)
$\frac{\partial q_{2,n}}{\partial t}(t,u)$ $=R_{n}(q_{1}(t,u),q_{2,n}(t,u)),\quad$ $q_{2,n}(0,u)=u_{2}\,.$ (2.14c)
We call $(P_{n},q_{1},q_{2,n})$ a solution to (2.14) if $P_{n}\in C^{1}([0,T],\mathbb{R})$, $q_{1}\in C^{1}([0,T],H)$, and $q_{2,n}\in C^{1}([0,T],\mathcal{S}_{n})$ solve (2.14a)–(2.14c).
Definition 2.5.
Denote by $\mathcal{S}_{n}^{\mathbb{C}}=\mathcal{S}_{n}\oplus\mathrm{i}\mathcal{S}^{n}$. Let $F_{n}$ be the analytic extension to $\mathcal{S}_{n}^{\mathbb{C}}$ of the function $F_{n}$ defined in (2.13) and $R_{n}$ be the analytic extension to $\mathbb{C}^{n}\times\mathcal{S}_{n}^{\mathbb{C}}$ of the function $R_{n}$ defined in (2.13). Consider the following system of ordinary differential equations
$\,\frac{\partial{\eta_{n}}}{\partial t}(t,u)$ $=F_{n}(\zeta_{2,n}(t,u)),\quad 0<t\leq T,\quad\eta_{n}(0,u)=0,$ (2.15a)
$\,\zeta_{1}(t,u)$ $=u_{1}+\mathcal{A}^{*}\int_{0}^{t}\zeta_{1}(s,u)\mathop{}\!\textnormal{d}s,\quad 0<t\leq T,\quad\zeta_{1}(0,u)=u_{1}\,$, (2.15b)
$\,\frac{\partial\zeta_{2,n}}{\partial t}(t,u)$ $=R_{n}(\zeta_{1}(t,u),\zeta_{2,n}(t,u)),\quad 0<t\leq T,\quad\zeta_{2,n}(0,u)=u_{2}.$ (2.15c)
For $u=(u_{1},u_{2})\in H^{\mathbb{C}}\times\mathcal{S}_{n}^{\mathbb{C}}$ and $T\geq 0$, we say that $(\eta_{n}(\cdot,u),\zeta_{1}(\cdot,u),\zeta_{2,n}(\cdot,u))\colon[0,T]\rightarrow\mathbb{R}\times H^{\mathbb{C}}\times\mathcal{S}_{n}^{\mathbb{C}}$ is a mild solution to (2.15a)-(2.15c) if $\eta_{n}(\cdot,u)\in C^{1}([0,T],\mathbb{R})$, $\zeta_{1}(\cdot,u)\in C^{1}([0,T],H^{\mathbb{C}})$ and $\zeta_{2,n}(\cdot,u)\in C^{1}([0,T],\mathcal{S}_{n}^{\mathbb{C}})$ satisfies (2.15a)-(2.15c). Then we call $(\eta_{n},\zeta_{1},\zeta_{2,n})$ a solution (up to time $T$ and with starting point $u$) of a complex Riccati system.
Next, we state an existence result for the complex Riccati system (2.15). The proof follows the same lines as Proposition 5.5; see also [32, Proposition 5.1] for the case $D=S_{d}^{+}$, $d\geq 2$.
Theorem 2.6.
Let $T>0$ and $n\in\mathbb{N}$. Let $Q,D\in\mathcal{S}_{1}^{+}(H)$, $\Upsilon\in H$, and $\mathbb{A}\in\mathcal{L}(H)$. Let $(\mathcal{A},\operatorname{dom}(\mathcal{A}))$ be the generator of a strongly continuous semigroup $(S_{t})_{t\geq 0}$ on $H$, and $(\mathcal{A}^{*},\operatorname{dom}(\mathcal{A}^{*}))$ its adjoint. Let $u=(u_{1},u_{2})\in\mathbb{C}^{n}\times(\mathcal{S}_{n}^{+})^{\mathbb{C}}$ and suppose that the real Riccati system (2.14a)–(2.14c) with initial data $\Re(u)=(\Re(u_{1}),\Re(u_{2}))$ admits a (unique) mild solution $(P_{n},q_{1,n},q_{2,n})$ on $[0,T]$, as guaranteed by Proposition 5.6. Then the complex Riccati system (2.15a)–(2.15c) with initial data $u$ admits a (necessarily unique) mild solution $(\eta_{n},\zeta_{1,n},\zeta_{2,n})$ on $[0,T]$.
2.2.4. Approximation of the Laplace transform of the infinite-dimensional Wishart stochastic volatility model
Our primary focus in this paper is the numerical implementation of option pricing in a setting where the forward price or rate follows an infinite-dimensional stochastic differential equation. To enable computations, we approximate the model by finite-dimensional systems.
For the Wishart stochastic volatility model, we do not perform a detailed convergence analysis of these approximations. Such an analysis, while important for theoretical completeness, would introduce significant technical complexity and is therefore left for future work. Our objective here is to demonstrate the practical feasibility and computational tractability of the method. For comparison, the case in which the covariance process is driven by jumps has been analyzed in [30, 31].
To keep the presentation concise and highlight the numerical methodology, we approximate the Laplace transform of the infinite-dimensional Wishart stochastic volatility model directly. Let $(X_{t},\mathcal{Y}_{t})_{t\geq 0}$ be as in (2.10), and let $(\eta_{n},\zeta_{1},\zeta_{2,n})$ be the solution to (2.15) with initial condition $(u_{1},u_{2})\in H^{\mathbb{C}}\times\mathcal{S}_{n}^{\mathbb{C}}$, whose existence is guaranteed by Proposition 2.6.
In our option pricing derivations in the next section, we use the following finite-rank approximation to the Laplace transform:
$\mathbb{E}\!\left[e^{\langle X_{T},u_{1}\rangle_{H^{\mathbb{C}}}-\operatorname{Tr}\big{(}\mathcal{Y}_{T}\,\iota_{n}u_{2}\iota_{n}^{*}\big{)}}\,\middle|\,\mathcal{F}_{t}\right]$
$\qquad\approx\exp\!\Big{(}-\eta_{n}(T-t,u)+\langle X_{t},\zeta_{1}(T-t,u)\rangle_{H}-\operatorname{Tr}\big{(}\mathcal{Y}_{t}\,\iota_{n}\zeta_{2,n}(T-t,u)\iota_{n}^{*}\big{)}\Big{)}.$ (2.16)
3. Pricing options on forwards in infinite-dimensional stochastic volatility models
We begin with a brief introduction to the modeling of commodity forward curves in the Heath-Jarrow-Morton framework, here we closely follow [4] and connect our two infinite-dimensional stochastic volatility models $(X_{t},\mathcal{Y}_{t})_{t\geq 0}$ and $(X_{t},{\sf Y}_{t})_{t\geq 0}$ to this framework.
Suppose $w\colon\mathbb{R}^{+}\rightarrow[1,\infty)$ is a non-decreasing function such that $w(0)=1$ and $\int_{0}^{\infty}w^{-1}(x)\,\mathop{}\!\textnormal{d}x<\infty$. Let $H_{w}$ be the space of absolutely continuous functions $f:\mathbb{R}^{+}\rightarrow\mathbb{R}$ such that
$\|f\|_{w}^{2}:=f(0)^{2}+\int_{0}^{\infty}w(x)|f^{\prime}(x)|^{2}\,\mathop{}\!\textnormal{d}x<\infty\,$,
where $f^{\prime}$ is the weak derivative of $f$. The space $H_{w}$ is a separable Hilbert space when endowed with the inner product
$\langle f,g\rangle_{w}:=f(0)g(0)+\int_{0}^{\infty}w(x)f^{\prime}(x)g^{\prime}(x)\,\mathop{}\!\textnormal{d}x.$
This space is proposed in [25] as a class of Hilbert spaces that match the economic reasoning about the forward rate curve in fixed income markets (see also [4] for the case of commodity forward markets).
The left-shift semigroup is a strongly continuous semigroup on $H_{w}$ with infinitesimal generator $\mathcal{A}$ being the operator of differentiation (in time-to-maturity coordinate). Moreover, the point evaluation $\delta_{x}(h)=u_{x}$ is a continuous linear functional on $H_{w}$, for all $x\in\mathbb{R}^{+}$. It can be expressed as (see [25, Lemma 5.3.1])
$\delta_{x}=\langle\cdot,u_{x}\rangle_{w}\,$, (3.1)
where for $x\in\mathbb{R}^{+}$, $u_{x}:\mathbb{R}^{+}\rightarrow\mathbb{R}$, $y\mapsto 1+\int_{0}^{x\wedge y}w^{-1}(z)\,\mathop{}\!\textnormal{d}z$.
Let $F(t,\tau)$ be the forward price at time $t>0$ for delivery at $\tau>t$. The log–forward in the Musiela parametrisation (time to maturity $x=\tau-t$) is
$f_{t}(x):=\ln F(t,t+x),\qquad x\in\mathbb{R}_{+}.$
Fix a weight function $w$ and assume the state process $(X_{t})_{t\geq 0}$ satisfies
$\mathop{}\!\textnormal{d}X_{t}=\big{(}\mathcal{A}X_{t}+C_{t}\big{)}\,\mathop{}\!\textnormal{d}t+D^{1/2}Y_{t}^{1/2}\,\mathop{}\!\textnormal{d}W_{t},\qquad X_{0}=x\in H_{w},$ (3.2)
where $\mathcal{A}=\partial_{x}$ is the derivative (shift) operator, $C$ is an integrable adapted $H_{w}$–valued process, $D\in\mathcal{L}(H_{w})$ is self-adjoint and positive, $(W_{t})_{t\geq 0}$ is a cylindrical Brownian motion, and $(Y_{t})_{t\geq 0}$ is either $(\mathcal{Y}_{t})_{t\geq 0}$ from (2.2.1) or $({\sf Y}_{t})_{t\geq 0}$ from (2.2). On a filtered probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{0\leq t\leq T},\mathbb{P})$, we model the (arbitrage-free) forward curve via
$F(t,t+x):=\exp\{\delta_{x}(X_{t})\}=\exp\{\langle X_{t},u_{x}\rangle_{w}\},\qquad 0\leq t\leq T,$ (3.3)
where $u_{x}$ is the Riesz representer of $\delta_{x}$ (see (3.1)). Assume $\xi$ is an $H_{w}$–valued adapted process such that
$\mathbb{E}\left[\exp\!\Big{\{}-\int_{0}^{T}\langle\xi_{s},\mathop{}\!\textnormal{d}W_{s}\rangle_{w}-\tfrac{1}{2}\int_{0}^{T}\|\xi_{s}\|_{w}^{2}\,\mathop{}\!\textnormal{d}s\Big{\}}\right]=1.$ (3.4)
By Girsanov’s theorem for cylindrical Brownian motion (e.g., [20, Thm. 13]),
$\widetilde{W}_{t}:=W_{t}-\int_{0}^{t}\xi_{s}\,\mathop{}\!\textnormal{d}s$
is a cylindrical Brownian motion under $\mathbb{Q}$ with density
$\left.\frac{\mathop{}\!\textnormal{d}\mathbb{Q}}{\mathop{}\!\textnormal{d}\mathbb{P}}\right|_{\mathcal{F}_{t}}=\exp\Big{\{}-\int_{0}^{t}\langle\xi_{s},\mathop{}\!\textnormal{d}W_{s}\rangle_{w}-\tfrac{1}{2}\int_{0}^{t}\|\xi_{s}\|_{w}^{2}\,\mathop{}\!\textnormal{d}s\Big{\}},\qquad 0\leq t\leq T.$
Under $\mathbb{Q}$, the process $(X_{t})_{t\geq 0}$ solves
$\mathop{}\!\textnormal{d}X_{t}=\big{(}\mathcal{A}X_{t}+C_{t}-D^{1/2}Y_{t}^{1/2}\xi_{t}\big{)}\,\mathop{}\!\textnormal{d}t+D^{1/2}Y_{t}^{1/2}\,\mathop{}\!\textnormal{d}\widetilde{W}_{t},\qquad 0\leq t\leq T.$ (3.5)
The following result gives a drift condition ensuring the martingale property of $t\mapsto F(t,\tau)$, for $t\leq\tau$. The proof follows the same lines as [7, Proposition 6.1]
Lemma 3.1.
. Let $\xi$ satisfy (3.4). Let $\mathcal{A}$, $C$, $D$, and $(Y_{t})_{t\geq 0}$ be as above, and let $S(t)$ be the left–shift semigroup generated by $\mathcal{A}$. Suppose there exists $h_{0}\in H_{w}$ such that, for all $0\leq t\leq\tau\leq T$,
$C_{t}-D^{1/2}Y_{t}^{1/2}\xi_{t}=-\tfrac{1}{2}\,D^{1/2}Y_{t}D^{1/2}\,u_{\tau-t}\qquad\text{with }u_{x}:=S^{*}(x)h_{0}.$ (3.6)
Then, for each fixed $\tau$, the process $t\mapsto F(t,\tau)=\exp\{\langle X_{t},u_{\tau-t}\rangle_{w}\}$ is a (local) $\mathbb{Q}$–martingale on $0,\tau]$.
From now on we impose the no–arbitrage drift condition (3.6). Equivalently, in the coupled SV dynamics (2.10) or (2.4) under $\mathbb{Q}$ we choose
$\Upsilon^{\vartheta}:=-\tfrac{1}{2}\,S^{*}(\vartheta)h_{0},\qquad\vartheta\geq 0,\ \ h_{0}\in H_{w}.$ (3.7)
Indeed, with this parametrization the model drift satisfies $C_{t}-D^{1/2}Y_{t}^{1/2}\xi_{t}=D^{1/2}Y_{t}D^{1/2}\Upsilon^{\vartheta}$, and for a claim with delivery lag $\vartheta=\tau-t$ this equals $\tfrac{1}{2}D^{1/2}Y_{t}D^{1/2}S^{*}(\tau-t)h_{0}$, which is exactly (3.6). Moreover, $\vartheta\mapsto D^{1/2}S^{*}(\vartheta)h_{0}\in H_{w}$ (see [7, Lem. 6.2]), so $\Upsilon^{\vartheta}\in H_{w}$ is well defined.
In pricing, all quantities depend on the delivery lag $\vartheta:=T_{1}-T_{0}$ through the evaluation functional $u_{\vartheta}=S^{*}(\vartheta)h_{0}$ and the choice $\Upsilon^{\vartheta}$ in (3.7). We therefore write the Riccati solutions as $(\eta^{\vartheta},\zeta^{\vartheta}_{1},\zeta^{\vartheta}_{2})$ for the Wishart specification and $(\phi^{\vartheta},\psi^{\vartheta}_{1},\psi^{\vartheta}_{2})$ for the pure-jump specification. For an option written at time $t$ on the forward delivering at $T_{1}$ and set at $T_{0}$ we have
$F(T_{0},T_{1})=\exp\{\langle X_{T_{0}},u_{T_{1}-T_{0}}\rangle_{w}\}=:F_{T_{0}}^{(\vartheta)},\qquad 0\leq T_{0}\leq T_{1}\leq T,$ (3.8)
a standard payoff in commodity markets such as oil, metals, and agriculture, see [6]. In the next proposition, we derive semi-explicit call prices via the Fourier transform of the payoff and the corresponding extended Riccati equations for our stochastic volatility models.
Proposition 3.2.
Let $T>0$, $0\leq T_{0}\leq T_{1}\leq T$, and $h_{0}\in H_{w}$. Set $\vartheta:=T_{1}-T_{0}$. Assume $(X^{\vartheta}_{t})_{t\geq 0}$ is as in (3.2), with $\Upsilon=-\tfrac{1}{2}S^{*}(\vartheta)h_{0}$, and let the forward price $F$ be defined by (3.8). Write $\mathbb{E}_{\mathbb{Q}}$ for expectation under $\mathbb{Q}$, and define $g:\mathbb{R}\to\mathbb{C}$ by
$g(\lambda):=\frac{K^{-(\nu-1+\mathrm{i}\lambda)}}{(\nu+\mathrm{i}\lambda)(\nu-1+\mathrm{i}\lambda)}\,,\qquad\nu>1,\ \lambda\in\mathbb{R}.$ (3.9)
Set $u:=(\nu+\mathrm{i}\lambda)\,u_{\vartheta}$. We consider the two covariance specifications below.
  1. (i)
    If the covariance process is $Y_{t}=\mathsf{Y}_{t}$ as in (2.2), and $(\phi^{\vartheta},\psi_{1}^{\vartheta},\psi_{2}^{\vartheta})$ denotes the solution to the generalized Riccati equations (2.8a)–(2.8c) on $[0,T]$ with initial data $(u,0)$, then for $0\leq t\leq T_{0}$ the time-$t$ price of the call with maturity $T_{0}$ and strike $K>0$ written on $F$ is
    $\mathbb{E}_{\mathbb{Q}}\big{[}(F(T_{0},T_{1})-K)^{+}\mid\mathcal{F}_{t}\big{]}=\frac{1}{2\pi}\int_{\mathbb{R}}g(\lambda)\,\exp\!\Big{(}$ $-\phi^{\vartheta}(T_{0}-t,u,0)+\langle\psi_{1}^{\vartheta}(T_{0}-t,u,0),X_{t}^{\vartheta}\rangle_{H_{w}^{\mathbb{C}}}$
    $\hskip 45.52458pt-\langle\psi_{2}^{\vartheta}(T_{0}-t,u,0),\mathsf{Y}_{t}\rangle_{\mathcal{H}^{\mathbb{C}}}\Big{)}\,d\lambda.$ (3.10)
  2. (ii)
    If the covariance process is $Y_{t}=\mathcal{Y}_{t}$ as in (2.2.1), and $(\eta^{\vartheta},\zeta_{1}^{\vartheta},\zeta_{2}^{\vartheta})$ denotes the solution to the Riccati equations (2.15a)–(2.15c) on $[0,T]$ with initial data $(u,0)$, then for $0\leq t\leq T_{0}$
    $\mathbb{E}_{\mathbb{Q}}\big{[}(F(T_{0},T_{1})-K)^{+}\mid\mathcal{F}_{t}\big{]}=\frac{1}{2\pi}\int_{\mathbb{R}}g(\lambda)\,\exp\!\Big{(}$ $-\eta^{\vartheta}(T_{0}-t,u,0)+\langle\zeta_{1}^{\vartheta}(T_{0}-t,u,0),X_{t}^{\vartheta}\rangle_{H_{w}^{\mathbb{C}}}$
    $\hskip 45.52458pt-\langle\zeta_{2}^{\vartheta}(T_{0}-t,u,0),\mathcal{Y}_{t}\rangle_{\mathcal{H}^{\mathbb{C}}}\Big{)}\,d\lambda.$ (3.11)
Proof.
Standard Fourier techniques for option pricing (see e.g., [23] or [26, Theorem 10.6]), yields
$\mathbb{E}_{\mathbb{Q}}[(F(T_{0},T_{1})-K)^{+}|\mathcal{F}_{t}]=\frac{1}{2\pi}\int_{\mathbb{R}}g(\lambda)\mathbb{E}_{\mathbb{Q}}[\operatorname{e}^{\langle X^{\vartheta}_{T_{0}},(\nu+\mathrm{i}\lambda)u_{\vartheta}\rangle_{H_{w}^{\mathbb{C}}}}|\mathcal{F}_{t}]\,\mathop{}\!\textnormal{d}\lambda\,.$
Then the statement follows from Theorems 2.3 and 2.6. ∎
Consider a put option with strike $K>0$ and exercise time $T_{0}$, i.e., the payoff is given by $(K-F(T_{0},T_{1}))^{+}$. Similarly to the case of a call option, we obtain the expressions in the right hand side of ((ii)) and ((i)) for the price of a put option respectively, written on $(X_{t}^{\vartheta},\mathcal{Y}_{t})_{t\geq 0}$ and $(X_{t}^{\vartheta},{\sf Y}_{t})_{t\geq 0}$ with $u=(\nu+\mathrm{i}\lambda)u_{\vartheta}$, for $\nu<0,\lambda\in\mathbb{R}$.
4. Numerical analysis
In this section we study several specifications of the coupled process $(X_{t},Y_{t})_{t\geq 0}$ and compute prices of options written on forwards. Our primary goal is to verify, across models, that the option values obtained from the affine, semi-explicit formulas ((i)) and ((ii)) match those produced by Monte Carlo (MC) simulation within sampling error and show that that the affine formulas allows for much more efficient pricing. We start by describing the general settings of our model specifications.
4.1. General settings
In the implementation, we choose as a weight function $w(x)=\operatorname{e}^{\alpha x}$, for some variable $\alpha>0$. The initial forward rate curve is chosen using a Nelson-Siegel curve. We know that $H_{w}$ is isometrically isomorphic to $\mathbb{R}\times L^{2}[0,\infty)$ via the isomorphism
$T(h)=(h(0),h^{\prime}w^{1/2})\,,\qquad T^{-1}(c,f)=c+\int_{0}^{\cdot}f(y)w^{-1/2}\,\mathop{}\!\textnormal{d}y.$ (4.1)
Assume $(e_{n})_{n\in N}$ is an orthonormal basis of $\mathbb{R}\times L^{2}[0,\infty)$, then $(f_{n})=(T^{-1}e_{n})_{n\in N}$ is an orthonormal basis of $H_{w}$. We choose the orthonormal basis $(1,0)$ and $(0,g_{n})$ for $\mathbb{R}\times L^{2}[0,\infty)$, where $(g_{n})_{n\in\mathbb{N}}$ are Laguerre polynomials multiplied by a weight function $\mathrm{e}^{-x/2}$ (see Appendix A). It follows that the orthogonal basis $(f_{n})_{n\in\mathbb{N}}$ for $H_{w}$ is given by
$f_{1}$ $=1\,$, (4.2)
$f_{n+1}$ $=\int_{0}^{\cdot}g_{n-1}(s)w^{-1/2}(s)\,\mathop{}\!\textnormal{d}s,\quad n\geq 1\,.$
We consider the option value at time $t=0$. We recall the representation of the evaluation operators $\delta_{x}$. It holds for any linear operator $\mathcal{V}\in\mathcal{L}(H_{w})$,
$\mathcal{V}^{*}f(x)=\delta_{x}\mathcal{V}^{*}f=\langle\mathcal{V}^{*}f,u_{x}\rangle_{w}=\langle f,\mathcal{V}u_{x}\rangle_{w}.$
We infer for $(S^{*}(t))_{t\geq 0}$ and the weight function $w(x)=\mathrm{e}^{\alpha x}$,
$S^{*}(t)h(x)$ $=h(0)+h(0)\int_{0}^{x\wedge t}w^{-1}(s)ds+\int_{t}^{x\vee t}\frac{h^{\prime}(s-t)w(s-t)}{w(s)}\mathop{}\!\textnormal{d}s$
$=h(0)+\frac{1}{\alpha}h(0)(1-\mathop{}\!\textnormal{e}^{-\alpha(x\wedge t)})+\mathop{}\!\textnormal{e}^{-\alpha t}\mathbbm{1}_{x\geq t}(h(x-t)-h(0))\,.$ (4.3)
We take the HJM drift $D^{1/2}\mathsf{Y}_{t}D^{1/2}\Upsilon$ in the drift of $(X_{t})_{t\geq 0}$ equal to $\frac{1}{2}D^{1/2}\mathsf{Y}_{t}D^{1/2}\mathcal{S}^{*}(\vartheta)h_{0}$ with $h_{0}\in H_{w}$. In the experiments on the stochastic volatility model with jumps, we consider two examples: the model driven by a Lévy subordinator with $B(Y_{t})=0$, and the so called Barndorff–Nielsen–Shephard (BNS) model, i.e., $B(\mathsf{Y}_{t})=(\mathbb{C}\mathsf{Y}_{t}+\mathsf{Y}_{t}\mathbb{C}^{*})$.
4.2. Validation of the MGF in the Infinite-Dimensional Wishart Stochastic Volatility Model
For the infinite-dimensional Wishart stochastic volatility model, we first validate the theoretical moment generating function (MGF) against its empirical counterpart obtained by Monte Carlo. This is the right diagnostic because, in light of (2.2.4), option prices are computed from a fixed finite-rank approximation; consistency of the MGF is therefore the primary check for pricing accuracy.
Throughout this experiment we work in a finite-rank setting: the Wishart process takes values in the cone of self-adjoint trace-class operators of rank at most $n$. We fix an arbitrary but finite $n\in\mathbb{N}$ and an orthonormal basis $(f_{k})_{k\in\mathbb{N}}$ of $H$ as introduced earlier, and project all operators onto $H_{n}:=\mathrm{span}\{f_{1},\dots,f_{n}\}$. The theoretical MGF is evaluated via the affine transform/Riccati solution on $H_{n}$, while the empirical MGF is computed from Monte Carlo samples as the sample average of $\exp\{\langle X_{T},u_{1}\rangle-\mathrm{Tr}(Y_{T}u_{2})\}$ for the same $(u_{1},u_{2})$.
We define the initial value $Y_{0}$ of the Wishart process as a rank-$n$ positive operator given by
$Y_{0}=\sum_{k=1}^{n}\lambda_{k}f_{k}\otimes f_{k},$
where $\lambda_{k}>0$ are constant eigenvalues chosen for numerical stability and interpretation. The volatility-driving Wishart process $(Y_{t})_{t\geq 0}$ is governed by the SDE
$\mathop{}\!\textnormal{d}Y_{t}=nQ\,\mathop{}\!\textnormal{d}t+Y_{t}\mathbb{A}\,\mathop{}\!\textnormal{d}t+\mathbb{A}^{*}Y_{t}\,\mathop{}\!\textnormal{d}t+\sqrt{Y_{t}}\,\mathop{}\!\textnormal{d}B_{t}\sqrt{Q}+\sqrt{Q}\,\mathop{}\!\textnormal{d}B_{t}^{*}\sqrt{Y_{t}},$
where $(B_{t})_{t\geq 0}$ is a cylindrical Brownian motion on $L^{2}(H)$, and $Q\in\mathbb{S}_{1}^{+}(H)$ and $\mathbb{A}\in\mathcal{L}(H)$ are model parameters. We specify the operator $Q$ as
$Q=\sum_{k=1}^{n}q_{k}f_{k}\otimes f_{k},$
with $q_{k}>0$ to control the volatility of the Wishart process along each principal direction $f_{k}$. The drift operator $\mathbb{A}$ is chosen to be diagonal in the basis $(f_{k})_{k\in\mathbb{N}}$, such that
$\mathbb{A}=-\sum_{k=1}^{n}a_{k}f_{k}\otimes f_{k},$
with $a_{k}>0$ ensuring mean reversion of the covariance process. In this example, we set the $q_{k}=a_{k}=\frac{1}{k^{2}}$, $k\in\mathbb{N}$. We use $D\in\mathbb{S}_{1}^{+}(H)$ for the volatility loading, defined as
$D=\frac{1}{2}\sum_{k=1}^{\infty}\frac{1}{k^{2}}f_{k}\otimes f_{k}.$
The Wishart process is coupled withththth the forward dynamics via the volatility-modulated SPDE given in equation (1.1).
A solution to the Riccati equations (2.15a)-(2.15c), with $F$ and $R$ specified in (2.11) and (2.12), is not available in closed form. In Appendix B, we present an efficient numerical scheme for solving these equations under the given parameterization. The method is flexible and can be adapted to alternative parameter choices.
Monte Carlo simulation of the Wishart model is considerably more demanding. The main difficulty lies in computing the square root of the operator-valued volatility process $Y_{t}$ at each time step, this is analytically intractable and must be approximated numerically. To limit computational overhead, we simulate only a single time step corresponding to a short horizon of $T_{0}=1/365$. However, at such a small horizon $T_{0}$, the inverse Fourier transform used for option pricing becomes numerically unstable. Therefore, instead of the option prices, we compare the moment generating function (MGF), which is the key intermediate quantity in affine models.
In this example, we find that the MGF computed from the affine representation aligns closely with the MGF estimated via Monte Carlo simulation, thereby confirming pricing consistency. This agreement is illustrated in Figure 1, where the Monte Carlo estimates converge to the affine-model MGF.
Figure 1. Moment generation function: Monte Carlo v.s. analytical formula, with $h_{0}=1,\alpha=0.1,T_{0}=1/365,N=5$.
4.3. Pricing in a pure-jump Lévy driven stochastic volatility model
As a second example we consider an affine pure-jump Lévy–type stochastic volatility specification in which the instantaneous covariance is an operator–valued Lévy subordinator.
Let $(f_{n})_{n\in\mathbb{N}}$ be the orthonormal basis of $H_{w}$ from (4.2), and define the self-adjoint, positive and diagonal in $(f_{n})_{n\in\mathbb{N}}$ operator
$D=\frac{1}{2}\sum_{n=1}^{\infty}\frac{1}{n^{2}}\,f_{n}\otimes f_{n}\in\mathcal{L}(H_{w}).$
Let $\mathcal{A}$ be the generator of the left-shift semigroup and $(\widetilde{W}_{t})_{t\geq 0}$ a cylindrical Brownian motion under $\mathbb{Q}$ as in Section 3. Consider a compound Poisson process $(J_{t})_{t\geq 0}$ with deterministic jump directions $D$ and intensity $\beta>0$:
$J_{t}=\sum_{i=1}^{N_{t}}D,\qquad N_{t}\sim\mathrm{Poisson}(\beta t).$
Let $\mathcal{H}_{w}$ denote the space of self-adjoint Hilbert–Schmidt operators on $H_{w}$ and $\mathcal{H}_{w}^{+}$ its positive cone. Under $\mathbb{Q}$ we take the coupled model
$\mathop{}\!\textnormal{d}X_{t}$ $=\big{(}\mathcal{A}X_{t}+D^{1/2}\mathsf{Y}_{t}D^{1/2}\Upsilon^{\vartheta}\big{)}\,\mathop{}\!\textnormal{d}t+D^{1/2}\mathsf{Y}_{t}^{1/2}\,\mathop{}\!\textnormal{d}\widetilde{W}_{t},\qquad X_{0}=x\in H_{w},$
$\mathop{}\!\textnormal{d}\mathsf{Y}_{t}$ $=\beta D\,\mathop{}\!\textnormal{d}t+\mathop{}\!\textnormal{d}J_{t},\qquad\mathsf{Y}_{0}=y\in\mathcal{H}_{w}^{+},$ (4.4)
with $\Upsilon^{\vartheta}=-\tfrac{1}{2}S^{*}(\vartheta)h_{0}$ as in (3.7).
The complex associated Riccati system (2.8a)–(2.8c) becomes
$\partial_{t}\phi(t,u)=F(\psi_{2}(t,u)),\qquad\psi_{1}(t,u)=u_{1}+\mathcal{A}^{*}\int_{0}^{t}\psi_{1}(s,u)\,\mathop{}\!\textnormal{d}s,\qquad\partial_{t}\psi_{2}(t,u)=R(\psi_{1}(t,u),\psi_{2}(t,u)),$
with
$R(h,u)=-\tfrac{1}{2}\,(D^{1/2}h)^{\otimes 2}-D^{1/2}h\otimes D^{1/2}\Upsilon^{\vartheta},\qquad F(u)=\beta\big{(}\mathop{}\!\textnormal{e}^{-\langle D,u\rangle}-1\big{)},$
The following lemma states closed-form expressions of the solutions to the system in the basis $(f_{n})$.
Lemma 4.1.
Let $w(x)=e^{\alpha x}$ with $\alpha>0$ and $(f_{n})_{n\in\mathbb{N}}$ as above. For $u_{1}=(\nu+i\lambda)\,u_{\vartheta}$ with $\nu>1$ and $\lambda\in\mathbb{R}$, the Riccati solution satisfies
$\psi_{1}(t,u)(x)$ $=(\nu+i\lambda)\!\left(1+\frac{1}{\alpha}\big{(}1-e^{-\alpha(x\wedge t)}\big{)}+\frac{1}{\alpha}e^{-\alpha t}\,\mathbf{1}_{\{x\geq t\}}\big{(}1-e^{-\alpha(\vartheta\wedge x-t)}\big{)}\right),$ (4.5)
$\psi_{2}(t,u)$ $=-\frac{(\nu+i\lambda)^{2}}{4}\sum_{i=1}^{\infty}\sum_{j=1}^{\infty}\frac{1}{ij}\,\langle f_{i},\cdot\rangle_{w}\,f_{j}\int_{0}^{t}f_{i}(s+\vartheta)f_{j}(s+\vartheta)\,\mathop{}\!\textnormal{d}s$
$\quad\;-\frac{(\nu+i\lambda)}{2}\sum_{i=1}^{\infty}\sum_{j=1}^{\infty}\frac{1}{ij}\,c_{j}(\vartheta)\,\langle f_{i},\cdot\rangle_{w}\,f_{j}\int_{0}^{t}f_{i}(s+\vartheta)\,\mathop{}\!\textnormal{d}s,$ (4.6)
where $D^{1/2}=\tfrac{1}{\sqrt{2}}\sum_{n\geq 1}\frac{1}{n}\,f_{n}\otimes f_{n}$ and $c_{j}(\vartheta):=\langle f_{j},S^{*}(\vartheta)h_{0}\rangle_{w}$, i.e., $D^{1/2}\Upsilon^{\vartheta}=\tfrac{1}{\sqrt{2}}\sum_{j\geq 1}\frac{1}{j}\,c_{j}(\vartheta)\,f_{j}$. Moreover,
$\langle D,\psi_{2}(t,u)\rangle$ $=-\frac{(\nu+i\lambda)^{2}}{8}\sum_{n=1}^{\infty}\frac{1}{n^{4}}\int_{0}^{t}f_{n}(s+\vartheta)^{2}\,\mathop{}\!\textnormal{d}s\;-\;\frac{(\nu+i\lambda)}{4}\sum_{n=1}^{\infty}\frac{c_{n}(\vartheta)}{n^{4}}\int_{0}^{t}f_{n}(s+\vartheta)\,\mathop{}\!\textnormal{d}s,$ (4.7)
and
$\phi(t,u)$ $=\beta t-\beta\int_{0}^{t}\exp{\left\{\frac{1}{8}(\nu+\mathrm{i}\lambda)^{2}\sum_{k=1}^{\infty}\frac{1}{k^{4}}\int_{0}^{s}f_{k}(l+\vartheta)\left(f_{k}(l+\vartheta)+\frac{2c_{k}(\vartheta)}{\nu+\mathrm{i}\lambda}\right)\,\mathop{}\!\textnormal{d}l\right\}}\,\mathop{}\!\textnormal{d}s\,.$ (4.8)
Proof.
By the semigroup property, $\psi_{1}(t,u)=S^{*}(t)u_{1}$, which gives (4.5). For the quadratic/linear drivers, write
$D^{1/2}\psi_{1}(s,u)=\frac{\nu+i\lambda}{\sqrt{2}}\sum_{n\geq 1}\frac{1}{n}\,f_{n}(s+\vartheta)\,f_{n},\qquad D^{1/2}\Upsilon^{\vartheta}=\frac{1}{\sqrt{2}}\sum_{j\geq 1}\frac{1}{j}\,c_{j}(\vartheta)\,f_{j}.$
Using $(a\otimes b)h=\langle a,h\rangle_{w}\,b$ and integrating
$\partial_{t}\psi_{2}=-\tfrac{1}{2}(D^{1/2}\psi_{1})^{\otimes 2}-D^{1/2}\psi_{1}\otimes D^{1/2}\Upsilon^{\vartheta}$ yields (4.6). For $\langle D,\psi_{2}\rangle=\sum_{n\geq 1}\langle Df_{n},\psi_{2}f_{n}\rangle_{w}$, use $Df_{n}=\tfrac{1}{2n^{2}}f_{n}$ and orthonormality to obtain (4.7). Finally, $\partial_{t}\phi=F(\psi_{2})$ with $F(u)=-\beta\big{(}e^{-\langle D,u\rangle}-1\big{)}$ gives (4.8). ∎
From Theorem 2.3 and Lemma 4.1 (with $u_{1}=(\nu+i\lambda)u_{\vartheta}$, $\nu>1$) we obtain
$\mathbb{E}_{\mathbb{Q}}\big{[}(F(t,t+\vartheta)-K)^{+}\mid\mathcal{F}_{0}\big{]}$ $=\frac{1}{2\pi}\int_{\mathbb{R}}g(\lambda)\,\exp\Big{(}-\phi(t,u)+\langle X_{0},\psi_{1}(t,u)\rangle_{w}-\langle Y_{0},\psi_{2}(t,u)\rangle\Big{)}\,\mathop{}\!\textnormal{d}\lambda$ (4.9)
$=\frac{1}{2\pi}\int_{\mathbb{R}}g(\lambda)\,\exp\Big{(}-\beta t+\beta\int_{0}^{t}\mathop{}\!\textnormal{e}^{-\langle D,\psi_{2}(s,u)\rangle}\,\mathop{}\!\textnormal{d}s$
$\qquad\qquad+(\nu+i\lambda)\,X_{0}(t+\vartheta)-\langle Y_{0},\psi_{2}(t,u)\rangle\Big{)}\,\mathop{}\!\textnormal{d}\lambda,$
where
$\langle D,\psi_{2}(s,u)\rangle=-\frac{(\nu+i\lambda)^{2}}{8}\sum_{n=1}^{\infty}\frac{1}{n^{4}}\int_{0}^{s}f_{n}(r+\vartheta)^{2}\,\mathop{}\!\textnormal{d}r-\frac{(\nu+i\lambda)}{4}\sum_{n=1}^{\infty}\frac{c_{n}(\vartheta)}{n^{4}}\int_{0}^{s}f_{n}(r+\vartheta)\,\mathop{}\!\textnormal{d}r,$
and $c_{n}(\vartheta):=\langle f_{n},S^{*}(\vartheta)h_{0}\rangle_{w}$.
For numerical implementation we truncate the series at $N$:
$\mathbb{E}_{\mathbb{Q}}\!\big{[}(F(t,t+\vartheta)-K)^{+}\mid\mathcal{F}_{0}\big{]}$ $\approx\frac{1}{2\pi}\int_{\mathbb{R}}g(\lambda)\,\exp\Big{(}-\beta t+\beta\int_{0}^{t}\mathop{}\!\textnormal{e}^{-\langle D,\psi_{2}^{(N)}(s,u)\rangle}\,\mathop{}\!\textnormal{d}s$ (4.10)
$\qquad\qquad\qquad\qquad+(\nu+i\lambda)X_{0}(t+\vartheta)-\langle Y_{0},\psi_{2}^{(N)}(t,u)\rangle\Big{)}\,\mathop{}\!\textnormal{d}\lambda,$
with
$\langle D,\psi_{2}^{(N)}(s,u)\rangle=-\frac{(\nu+i\lambda)^{2}}{8}\sum_{n=1}^{N}\frac{1}{n^{4}}\int_{0}^{s}f_{n}(r+\vartheta)^{2}\,\mathop{}\!\textnormal{d}r-\frac{(\nu+i\lambda)}{4}\sum_{n=1}^{N}\frac{c_{n}(\vartheta)}{n^{4}}\int_{0}^{s}f_{n}(r+\vartheta)\,\mathop{}\!\textnormal{d}r.$
The function $f_{n}$ can be efficiently evaluated by iteration, see Appendix A, and consequently the integrals $\int_{0}^{s}f_{n}(r+\vartheta)\,\mathop{}\!\textnormal{d}r$ and $\int_{0}^{s}f_{n}(r+\vartheta)^{2}\,\mathop{}\!\textnormal{d}r$ can be calculated by using Gaussian quadrature method.
Tables 1 and 2 present the convergence analysis on the number of basis functions used, i.e., $N$. We use $N=10$ as the benchmark, and compute the relative absolute differences in the option price for $N=2,3,5,8$ compared to $N=10$, with different time to maturity $\vartheta$ (in Table 1) and different intensity $\beta$ (in Table 2). The results show that the difference is already less than 0.02% when $N=5$, which suggest the option price converges fast w.r.t the number of used basis function. Moreover, Figure 2 shows the comparison on the option prices between the Monte Carlo simulation and the analytical formula. One easily sees that the forward curve generated by Monte Carlo simulation converges to the curve calculated by our analytical formula.
Time to maturity 1 3 5 10 20
N=2 0.06% 0.09% 0.33% 0.63% 0.65%
N=3 0.01% 0.04% 0.03% 0.05% 0.04%
N=5 0.01% 0.02% 0.01% 0.00% 0.00%
N=8 0.00% 0.00% 0.00% 0.00% 0.00%
Table 1. Relative absolute difference for different number of basis function $n$ and time to maturity $\vartheta$ on the option price compared to $n=10$, with $Y_{0}=D,h_{0}=1,\alpha=0.1,T_{0}=1$, strike $K=1$ and the intensity $\beta=1$. Pure jump case.
beta 0.5 1 2 5
N=2 1.46% 1.26% 1.01% 0.60%
N=3 0.10% 0.05% 0.02% 0.02%
N=5 0.01% 0.01% 0.01% 0.01%
N=8 0.01% 0.01% 0.01% 0.00%
Table 2. Relative absolute difference for different $n$ and $\beta$ on the option price compared to $n=10$, with $Y_{0}=D,h_{0}=1,\alpha=0.1,T_{0}=2$, strike $K=2$ and the time to maturity $\vartheta=1$. Pure jump case.
Figure 2. Monte Carlo v.s. analytical formula, with $Y_{0}=D,h_{0}=1,\alpha=0.1,T_{0}=1,N=5$, strike $K=1$ and the intensity $\beta=1$.
4.4. Pricing in the operator-valued Barndorff–Nielsen–Shephard stochastic volatility model
Let $\mathcal{A}$, $D$, $(\widetilde{W}_{t})_{t\geq 0}$, and $(J_{t})_{t\geq 0}$ be as in Example 4.3, with $J_{t}=\sum_{i=1}^{N_{t}}D$ a compound Poisson process of intensity $\beta>0$, independent of $\widetilde{W}$. Let $\mathcal{H}_{w}$ be the space of self-adjoint Hilbert–Schmidt operators on $H_{w}$ and $\mathcal{H}_{w}^{+}$ its positive cone. We consider the operator-valued BNS dynamics
$\mathop{}\!\textnormal{d}X_{t}$ $=\big{(}\mathcal{A}X_{t}+D^{1/2}\mathsf{Y}_{t}D^{1/2}\Upsilon^{\vartheta}\big{)}\,\mathop{}\!\textnormal{d}t+D^{1/2}\mathsf{Y}_{t}^{1/2}\,\mathop{}\!\textnormal{d}\widetilde{W}_{t},\qquad X_{0}=x\in H_{w},$
$\mathop{}\!\textnormal{d}\mathsf{Y}_{t}$ $=\big{(}Q\mathsf{Y}_{t}+\mathsf{Y}_{t}Q^{*}\big{)}\,\mathop{}\!\textnormal{d}t+\mathop{}\!\textnormal{d}J_{t},\qquad\mathsf{Y}_{0}=y\in\mathcal{H}_{w}^{+},$ (4.11)
where $Q\in\mathcal{L}(H_{w})$ is a (mean-reverting) drift operator (we will take it diagonal below) and $\Upsilon^{\vartheta}=-\tfrac{1}{2}S^{*}(\vartheta)h_{0}$ as in (3.7). The associated complex Riccati system (2.8a)–(2.8c) becomes
$\partial_{t}\phi(t,u)=F(\psi_{2}(t,u)),\qquad\psi_{1}(t,u)=u_{1}+\mathcal{A}^{*}\int_{0}^{t}\psi_{1}(s,u)\,\mathop{}\!\textnormal{d}s,\qquad\partial_{t}\psi_{2}(t,u)=R(\psi_{1}(t,u),\psi_{2}(t,u)),$
with maps
$R(h,u)=uQ^{*}+Qu-\tfrac{1}{2}\,(D^{1/2}h)^{\otimes 2}-D^{1/2}h\otimes D^{1/2}\Upsilon^{\vartheta},\qquad F(u)=\beta\big{(}\mathop{}\!\textnormal{e}^{-\langle D,u\rangle}-1\big{)},$
where $\langle\cdot,\cdot\rangle$ denotes the Hilbert–Schmidt pairing on $\mathcal{H}_{w}$.
Lemma 4.2.
Let $w(x)=e^{\alpha x}$ with $\alpha>0$ and $(f_{n})_{n\in\mathbb{N}}$ be the orthonormal basis from (4.2). Assume the OU drift is diagonal,
$Q\;=\;-\sum_{n\geq 1}a_{n}\,f_{n}\otimes f_{n},\qquad a_{n}\geq 0,$
and set $D^{1/2}=\tfrac{1}{\sqrt{2}}\sum_{n\geq 1}\tfrac{1}{n}\,f_{n}\otimes f_{n}$, $c_{j}(\vartheta):=\langle f_{j},S^{*}(\vartheta)h_{0}\rangle_{w}$ (so $D^{1/2}\Upsilon^{\vartheta}=\tfrac{1}{\sqrt{2}}\sum_{j\geq 1}\tfrac{1}{j}c_{j}(\vartheta)\,f_{j}$). For $u_{1}=(\nu+i\lambda)u_{\vartheta}$ with $\nu>1$, $\lambda\in\mathbb{R}$, the complex Riccati solution to (4.4) satisfies
$\psi_{1}(t,u)(x)$ $=(\nu+i\lambda)\!\left(1+\frac{1}{\alpha}\big{(}1-e^{-\alpha(x\wedge t)}\big{)}+\frac{1}{\alpha}e^{-\alpha t}\,\mathbf{1}_{\{x\geq t\}}\big{(}1-\mathop{}\!\textnormal{e}^{-\alpha(\vartheta\wedge x-t)}\big{)}\right),$ (4.12)
$\psi_{2}(t,u)f_{n}$ $=-\sum_{j\geq 1}\int_{0}^{t}\mathop{}\!\textnormal{e}^{-(a_{n}+a_{j})(t-s)}\Big{[}\frac{(\nu+i\lambda)^{2}}{4\,nj}\,f_{n}(s+\vartheta)f_{j}(s+\vartheta)+\frac{(\nu+i\lambda)}{2\,nj}\,c_{j}(\vartheta)\,f_{n}(s+\vartheta)\Big{]}\mathop{}\!\textnormal{d}sf_{j},$ (4.13)
and
$\phi(t,u)$ $=\beta t-\beta\int_{0}^{t}\exp\big{(}\frac{1}{8}(\nu+\mathrm{i}\lambda)^{2}\sum_{n=1}^{\infty}\frac{1}{n^{4}}e^{\frac{s}{n^{2}}}\int_{0}^{s}\mathop{}\!\textnormal{e}^{-\frac{l}{n^{2}}}f_{n}(l+\vartheta)\left(f_{n}(l+\vartheta)+\frac{2c_{n}(\vartheta)}{\nu+\mathrm{i}\lambda}\right)\,\mathop{}\!\textnormal{d}l\big{)}\,\mathop{}\!\textnormal{d}s.$ (4.14)
In particular, we have
$\langle D,\psi_{2}(t,u)\rangle=-\int_{0}^{t}\Big{[}\frac{(\nu+i\lambda)^{2}}{8}\sum_{n\geq 1}\frac{1}{n^{4}}f_{n}(s+\vartheta)^{2}+\frac{(\nu+i\lambda)}{4}\sum_{n\geq 1}\frac{c_{n}(\vartheta)}{n^{4}}f_{n}(s+\vartheta)\Big{]}\mathop{}\!\textnormal{e}^{-2a_{n}(t-s)}\,\mathop{}\!\textnormal{d}s.$ (4.15)
Proof.
Since $\psi_{1}$ solves $\psi_{1}(t,u)=u_{1}+\mathcal{A}^{*}\int_{0}^{t}\psi_{1}(s,u)\,\mathop{}\!\textnormal{d}s$ with $u_{1}=(\nu+i\lambda)u_{\vartheta}$, we have the mild form $\psi_{1}(t,u)=S^{*}(t)u_{1}$, which yields (4.12) exactly as in Lemma 4.1.
The $\psi_{2}$-equation reads
$\partial_{t}\psi_{2}(t)=\underbrace{\psi_{2}(t)Q^{*}+Q\psi_{2}(t)}_{\mathcal{L}_{Q}(\psi_{2}(t))}\;-\;\frac{1}{2}\,(D^{1/2}\psi_{1}(t))^{\otimes 2}\;-\;D^{1/2}\psi_{1}(t)\otimes D^{1/2}\Upsilon^{\vartheta},\qquad\psi_{2}(0)=0.$
Hence, by variation of constants,
$\psi_{2}(t)=\int_{0}^{t}\mathop{}\!\textnormal{e}^{(t-s)\mathcal{L}_{Q}}\left[\,-\frac{1}{2}(D^{1/2}\psi_{1}(s))^{\otimes 2}-D^{1/2}\psi_{1}(s)\otimes D^{1/2}\Upsilon^{\vartheta}\right]\mathop{}\!\textnormal{d}s.$ (4.16)
Assume $Q=-\sum_{n\geq 1}a_{n}\,f_{n}\otimes f_{n}$ (with $a_{n}\geq 0$). Then, for basis tensors,
$\mathop{}\!\textnormal{e}^{(t-s)\mathcal{L}_{Q}}(f_{n}\otimes f_{j})=\mathop{}\!\textnormal{e}^{-(a_{n}+a_{j})(t-s)}\,f_{n}\otimes f_{j}.$
Moreover,
$D^{1/2}\psi_{1}(s)=\frac{\nu+i\lambda}{\sqrt{2}}\sum_{n\geq 1}\frac{1}{n}\,f_{n}(s+\vartheta)\,f_{n},\qquad D^{1/2}\Upsilon^{\vartheta}\;=\;\frac{1}{\sqrt{2}}\sum_{j\geq 1}\frac{1}{j}\,c_{j}(\vartheta)\,f_{j},\quad c_{j}(\vartheta):=\langle f_{j},S^{*}(\vartheta)h_{0}\rangle_{w}.$
Using $(a\otimes b)h=\langle a,h\rangle_{w}\,b$ and (4.16), we obtain for each $n$:
$\psi_{2}(t)f_{n}$ $=-\!\sum_{j\geq 1}\!\int_{0}^{t}e^{-(a_{n}+a_{j})(t-s)}\Big{[}\frac{(\nu+i\lambda)^{2}}{4\,nj}\,f_{n}(s+\vartheta)f_{j}(s+\vartheta)+\frac{(\nu+i\lambda)}{2\,nj}\,c_{j}(\vartheta)\,f_{n}(s+\vartheta)\Big{]}\mathop{}\!\textnormal{d}s\;f_{j},$
which is precisely (4.13). Since $D=\frac{1}{2}\sum_{n\geq 1}\frac{1}{n^{2}}\,f_{n}\otimes f_{n}$,
$\langle D,\psi_{2}(t)\rangle=\sum_{n\geq 1}\frac{1}{2n^{2}}\,\langle f_{n},\psi_{2}(t)f_{n}\rangle_{w}.$
By orthonormality, only the $j=n$ terms survive in $\langle f_{n},\psi_{2}(t)f_{n}\rangle_{w}$, giving
$\langle f_{n},\psi_{2}(t)f_{n}\rangle_{w}=-\int_{0}^{t}e^{-2a_{n}(t-s)}\left[\frac{(\nu+i\lambda)^{2}}{4\,n^{2}}\,f_{n}(s+\vartheta)^{2}+\frac{(\nu+i\lambda)}{2\,n^{2}}\,c_{n}(\vartheta)\,f_{n}(s+\vartheta)\right]\mathop{}\!\textnormal{d}s,$
from which we deduce formula (4.15). Finally, since for the compound-Poisson input without finite-variation drift it holds that $\partial_{t}\phi(t,u)=F(\psi_{2}(t,u))=\beta\big{(}e^{-\langle D,\psi_{2}(t,u)\rangle}-1\big{)}$, integrating over $[0,t]$ yields (4.14). ∎
Therefore, the option price is
$\mathbb{E}_{\mathbb{Q}}\big{[}(F(t,t+\vartheta)-K)^{+}\mid\mathcal{F}_{0}\big{]}$ $=\frac{1}{2\pi}\int_{\mathbb{R}}g(\lambda)\,\exp\!\Big{(}-\phi(t,u)+\langle X_{0},\psi_{1}(t,u)\rangle_{w}-\langle Y_{0},\psi_{2}(t,u)\rangle\Big{)}\,\mathop{}\!\textnormal{d}\lambda$ (4.17)
$=\frac{1}{2\pi}\int_{\mathbb{R}}g(\lambda)\,\exp\!\Big{(}-\beta\!\int_{0}^{t}\!\big{(}\mathop{}\!\textnormal{e}^{-\langle D,\psi_{2}(s,u)\rangle}-1\big{)}\,\mathop{}\!\textnormal{d}s\big{)}$
$\qquad\qquad\times\exp\!\Big{(}(\nu+i\lambda)\,X_{0}(t+\vartheta)-\langle Y_{0},\psi_{2}(t,u)\rangle\Big{)}\,\mathop{}\!\textnormal{d}\lambda,$
where $u_{1}=(\nu+i\lambda)u_{\vartheta}$ with $\nu>1$, and $\langle D,\psi_{2}(t,u)\rangle$ is as in (4.15), with $Q=-\sum_{n\geq 1}a_{n}\,f_{n}\otimes f_{n}$ ($a_{n}\geq 0$) and $c_{n}(\vartheta):=\langle f_{n},S^{*}(\vartheta)h_{0}\rangle_{w}$. The term $\langle X_{0},\psi_{1}(t,u)\rangle_{w}=(\nu+i\lambda)X_{0}(t+\vartheta)$ comes from $\psi_{1}(t,u)=S^{*}(t)u_{1}$.
For the numerics, we truncate the series at $N$:
$\mathbb{E}_{\mathbb{Q}}\big{[}(F(t,t+\vartheta)-K)^{+}\mid\mathcal{F}_{0}\big{]}$ $\approx\frac{1}{2\pi}\int_{\mathbb{R}}g(\lambda)\,\exp\Big{(}-\beta\!\int_{0}^{t}\!\big{(}e^{-\langle D,\psi_{2}^{(N)}(s,u)\rangle}-1\big{)}\,\mathop{}\!\textnormal{d}s\Big{)}$ (4.18)
$\qquad\qquad\times\exp\Big{(}(\nu+i\lambda)\,X_{0}(t+\vartheta)-\langle Y_{0},\psi_{2}^{(N)}(t,u)\rangle\Big{)}\,\mathop{}\!\textnormal{d}\lambda,$
with
$\langle D,\psi_{2}^{(N)}(t,u)\rangle=-\int_{0}^{t}\Big{[}\frac{(\nu+i\lambda)^{2}}{8}\sum_{n=1}^{N}\frac{1}{n^{4}}f_{n}(s+\vartheta)^{2}+\frac{(\nu+i\lambda)}{4}\sum_{n=1}^{N}\frac{c_{n}(\vartheta)}{n^{4}}f_{n}(s+\vartheta)\Big{]}\mathop{}\!\textnormal{e}^{-2a_{n}(t-s)}\,\mathop{}\!\textnormal{d}s.$
In our experiments, we specialize to $a_{n}=\frac{1}{2n^{2}}$, then $e^{-2a_{n}(t-s)}=\mathop{}\!\textnormal{e}^{-(t-s)/n^{2}}$.
The following Figure 3 and Tables 3 and 4 present the Monte Carlo convergence and the convergence analysis of number of used basis function $N$. Similar to the results in the pure jump case, one observes the option price converges fast w.r.t the number of used basis function, and Monte Carlo simulation converges to our analytical formula.
Time to maturity 1 3 5 10 20
N=2 0.03% 0.07% 0.23% 0.42% 0.42%
N=3 0.01% 0.03% 0.02% 0.03% 0.01%
N=5 0.01% 0.02% 0.00% 0.00% 0.00%
N=8 0.00% 0.00% 0.00% 0.00% 0.00%
Table 3. Relative absolute difference for different number of basis function $N$ and time to maturity $\vartheta$ on the option price compared to $N=10$, with $Y_{0}=D,h_{0}=1,\alpha=0.1,T_{0}=1$, strike $K=1$ and the intensity $\beta=1$. BNS model
beta 0.5 1 2 5
N=2 0.37% 0.32% 0.24% 0.03%
N=3 0.03% 0.05% 0.10% 0.21%
N=5 0.00% 0.00% 0.00% 0.00%
N=8 0.00% 0.00% 0.00% 0.00%
Table 4. Relative absolute difference for different NN and $\beta$ on the option price compared to $N=10$, with $Y_{0}=D,h_{0}=1,\alpha=0.1,T_{0}=2$, strike $K=2$ and the time to maturity $\vartheta=1$. BNS model.
Figure 3. Monte Carlo v.s. analytical formula, with $Y_{0}=D,h_{0}=1,\alpha=0.1,T_{0}=1,N=5$, strike $K=1$ and the intensity $\beta=1$.
4.5. Running time analysis
Table 5 presents the runtime comparison for option pricing between the affine semi-analytical method and the Monte Carlo simulation, based on the experiments (Monte Carlo v.s. analytical formula) in Section 4.3 and 4.4. Both experiments were conducted on a machine equipped with an AMD Ryzen 5 PRO 7545U CPU (6 cores, 3.20 GHz base clock), 32 GB DDR5 RAM. Algorithms were implemented in Python 3.12.8 using NumPy 1.26.4 and SciPy 1.14.1. Each reported runtime is the average over 10 independent runs, measured with the "time" module. The results clearly show that the affine semi-analytical method is approximately 6000 times faster than the Monte Carlo simulation, which indicates its significant advantage for practical applications where computational efficiency is critical.
Runtime (seconds) Affine formula Monte Carlo
Lévy driven model 0.03 173.70
BNS model 0.03 181.79
Table 5. Runtime (seconds) comparison: Affine formula v.s. Monte Carlo (1 million scenarios). On the option pricing for pure-jump Lévy driven model (Section 4.3) and the BNS model (Section 4.4)
5. Computation of complex moments
In this section, we prove the existence of solutions to the complex-extensions of the generalized Riccati equations (2.8a)-(2.8c) and (2.15a)-(2.15c). These existence results are then used to compute exponential moments for the affine pure-jump stochastic volatility model and to derive an approximation for the Laplace transform of the Wishart stochastic volatility model. To this end, we introduce the concept of quasi-monotonicity.
Definition 5.1.
Let $V$ be a real Banach space, $K\subset V$ a wedge. Given $D\subset V$ and $f\colon D\to V$ is said to be quasi monotone (with respect to $K$) if $\forall x,y\in D$ satisfying $x\leq_{K}y$ and $\forall x^{\prime}\in K^{*}$ satisfying $\langle x^{\prime},x-y\rangle_{V^{\prime},V}=0$, it holds $\langle x^{\prime},f(x)-f(y)\rangle_{V^{\prime},V}\leq 0\,.$
5.1. Real and complex moments for the affine pure-jump stochastic volatility model
In the following proposition we show the existence of a unique solution to the extended Riccati equations (2.7a)-(2.7c) up to its maximal lifetime.
Proposition 5.2.
Let $(b,B,m,\mu)$ be an admissible parameter set satisfying Assumptions Assumption $\mathfrak{A}$ and let $(\mathcal{A},\operatorname{dom}(\mathcal{A}))$ be the generator of a strongly continuous semigroup, let $D\in\mathcal{L}(H)$ be a positive self-adjoint operator, and let $\Upsilon\in H$. Then for every $u=(u_{1},u_{2})\in H\times\mathcal{O}$ there exist a positive real number $T_{\sf q_{2}}$ and a unique solution $({\sf p}(\cdot,u),{\sf q}_{1}(\cdot,u),+{\sf q}_{2}(\cdot,u))$ to (2.7a)-(2.7c) on $0,T_{\sf q_{2}})$.
Proof.
Standard semigroup theory ensures that for any $u_{1}\in H$ and $T>0$ the unique mild solution to (2.7b) is given by ${\sf q}_{1}(t,u)=S^{*}(t)u_{1}$ for $t\in[0,T]$. We define $\mathcal{R}_{u_{1}}(t,\cdot)\colon\mathcal{O}\to\mathcal{H}$, by
$\mathcal{R}_{u_{1}}(t,u)$ $=\tilde{B}^{*}(u)-\tfrac{1}{2}\big{(}D^{1/2}S^{*}(t)u_{1}\big{)}^{\otimes 2}-D^{1/2}S^{*}(t)u_{1}\otimes D^{1/2}\Upsilon$
$\qquad-\int_{\mathcal{H}^{+}\setminus\{0\}}\big{(}\mathop{}\!\textnormal{e}^{-\langle\xi,u\rangle}-1\big{)}\frac{\mu(\mathop{}\!\textnormal{d}\xi)}{\|\xi\|_{2}^{2}}.$
Plugging ${\sf q}_{1}(t,u)$ into (2.7c), we thus obtain the equation
$\begin{cases}\frac{\partial{\sf q}_{2}}{\partial t}(t,u)=\mathcal{R}_{u_{1}}(t,{\sf q}_{2}(t,u))\,,\\ {\sf q}_{2}(0,u)=u_{2}\,.\end{cases}$ (5.1)
Observe that the function $\mathcal{R}_{u_{1}}(t,\cdot)$ is locally Lipschitz continuous on $\mathcal{O}$ for every $t\geq 0$ and $u_{1}\in H$. Since $\mathcal{O}$ is assumed to be open, it follows from standard ODE results (see e.g. [36, Chapter 6, Proposition 1.2]) that for every $u_{2}\in\mathcal{O}$ there exist a $T_{\sf q_{2}}>0$ and a unique solution ${\sf q}_{2}(\cdot,u)$ of (5.1) on $0,T_{{\sf q}_{2}})$ with
$T_{\sf q_{2}}=\liminf_{n\to\infty}\left\{t\geq 0\colon\|{\sf q}_{2}(t,u)\|_{2}\geq n\text{ or }{\sf q}_{2}(t,u)\in\partial\mathcal{O}\right\}.$
By inserting $\sf q_{2}(\cdot,u)$ in (2.7a) and observing that $F$ is continuous on $\mathcal{O}$, the statement follows. ∎
To compute exponential moments for the process $(X_{t},\mathsf{Y}_{t})_{t\geq 0}$ in (2.4), we first consider the joint process $(X^{(n)}_{t},\mathsf{Y}_{t})_{t\geq 0}$ obtained by replacing $\mathcal{A}$ in (2.4) by its Yosida approximation $\mathcal{A}^{(n)}:=n\mathcal{A}(nI-\mathcal{A})^{-1}$. Namely, we consider the process $X^{(n)}\colon[0,\infty)\times\Omega\rightarrow H$ given by the solution to (see [21, Proposition 6.4] for the existence of the solution to $X^{(n)}$)
$X^{(n)}_{t}=x+\int_{0}^{t}(\mathcal{A}^{(n)}X^{(n)}_{s}+D^{1/2}X^{(n)}_{s}D^{1/2}\Upsilon)\,\mathop{}\!\textnormal{d}s+\int_{0}^{t}D^{1/2}X_{s}^{1/2}\,\mathop{}\!\textnormal{d}W_{s}\,,\quad t\geq 0.$ (5.2)
The use of the approximation allows us to exploit the semimartingale theory and to apply the Itô formula and standard techniques in order to compute exponential moments for $(X^{(n)}_{t},\mathsf{Y}_{t})_{t\geq 0}$. Then we take the limit when $n$ goes to $\infty$. Observe that from [21, Proposition 7.5] it holds
$\lim_{n\rightarrow\infty}\mathbb{E}\left[\sup_{0\leq t\leq T}\|X^{(n)}_{t}-X_{t}\|^{2}_{H}\right]=0\,.$ (5.3)
We are ready to derive the exponential moments of the affine pure-jump stochastic volatility model.
Proposition 5.3.
Let $(b,B,m,\mu)$ be an admissible parameter set satisfying Assumptions Assumption $\mathfrak{A}$ and Assumption $\mathfrak{B}$, let $(X,Y)$ be the stochastic volatility model (2.4), let $u=(u_{1},u_{2})\in H\times\mathcal{O}$, and let $({\sf P}(\cdot,u),{\sf q}_{1}(\cdot),u),{\sf q}_{2}(\cdot,u))$ be the mild solution to the Riccati equation (2.7) up to time $T$ the existence of which is guaranteed by Proposition 5.2. Then we have:
$\mathbb{E}\left[\mathop{}\!\textnormal{e}^{\langle X_{T},u_{1}\rangle_{H}-\langle{\sf Y}_{T},u_{2}\rangle}\right]<\infty.$ (5.4)
Moreover, for every $t\leq T$, it holds
$\mathbb{E}\left[\mathop{}\!\textnormal{e}^{\langle X_{T},u_{1}\rangle_{H}-\langle{\sf Y}_{T},u_{2}\rangle}\mid\mathcal{F}_{t}\right]=\mathop{}\!\textnormal{e}^{-{\sf P}(T-t,u)+\langle X_{t},{\sf q}_{1}(T-t,u)\rangle_{H}-\langle{\sf Y}_{t},{\sf q}_{2}(T-t,u)\rangle}.$ (5.5)
Proof.
Let $({\sf P}^{(n)}(\cdot,u),{\sf q}_{1}^{(n)}(\cdot,u),{\sf q}_{2}^{(n)}(\cdot,u))$ be the solution to (2.7) with $\mathcal{A}=\mathcal{A}^{(n)}$ (the $n$th Yosida approximation). We define the function $g_{u}^{(n)}(t,y,x)\colon[0,T]\times H\times\mathcal{H}^{+}\rightarrow\mathbb{R}$ as follows
$g_{u}^{(n)}(t,x,y)=\exp{\{-{\sf P}^{(n)}(T-t,u)+\langle x,{\sf q}^{(n)}_{1}(T-t,u)\rangle_{H}-\langle y,{\sf q}^{(n)}_{2}(T-t,u)\rangle\}}\,.$
Similarly to the proof of [16, Theorem 3.3], we apply the Itô’s formula to $g_{u}^{(n)}(t,X^{(n)}_{t},{\sf Y}_{t})$ to deduce that it is a local martingale.
Then observing that $g_{u}^{(n)}(t,X_{t}^{(n)},{\sf Y}_{t})$ is strictly positive for all $t\in[0,T]$, we infer it is a supermartingale and we have
$\mathbb{E}\left[\mathop{}\!\textnormal{e}^{\langle X^{(n)}_{T},u\rangle_{H}-\langle{\sf Y}_{T},u\rangle}\right]$ $=\mathbb{E}\left[g_{u}^{(n)}(T,X_{T}^{(n)},{\sf Y}_{T})\right]$
$\leq g_{u}^{(n)}(0,x,y)<\infty\,,\quad(x,y)\in H\times\mathcal{H}\,$,
from which it follows that $(g_{u}^{(n)}(t,X_{t}^{(n)},{\sf Y}_{t}))_{t\in[0,T]}$ is a martingale and hence it holds
$\mathbb{E}\left[\mathop{}\!\textnormal{e}^{\langle X_{T}^{(n)},u_{1}\rangle_{H}-\langle{\sf Y}_{T},u_{2}\rangle}\mid\mathcal{F}_{t}\right]$ $=\mathop{}\!\textnormal{e}^{-{\sf P}^{(n)}(T-t,u)+\langle X_{t},{\sf q}_{1}^{(n)}(T-t,u)\rangle_{H}-\langle{\sf Y}_{t},{\sf q}_{2}^{(n)}(T-t,u)\rangle}.$
Following similar arguments as in the proof of [16, Proposition 3.5], we deduce that
$\lim_{n\rightarrow\infty}\sup_{t\in[0,T]}|{\sf P}^{(n)}(t,u)-{\sf P}(t,u)|=0$
and
$\lim_{n\rightarrow\infty}\left(\sup_{t\in[0,T]}\|{\sf q}_{1}^{(n)}(t,u)-{\sf q}_{1}(t,u)\|_{H}+\sup_{t\in[0,T]}\|{\sf q}_{2}^{(n)}(t,u)-{\sf q}_{2}(t,u)\|_{2}\right)\,.$
Hence taking limits for $n\rightarrow\infty$ and invoking (5.3) yields the statement. ∎
Before proving the existence and uniqueness of a solution to (2.8a)-(2.8c), we give the following lemma:
Lemma 5.4.
Let $\sf R$} be as in (2.6). Then there exists a locally Lipschitz continuous function $g$ on $\mathcal{H}$ such that for all $h\in H^{\mathbb{C}}$ and $u\in S(\mathcal{O})$ we have
$\Re(\langle\overline{u},{\sf R}(h,u)\rangle)\leq g(\Re(u))(1+\|u\|_{\mathcal{H}^{\mathbb{C}}}^{2})(1+\|h\|_{H^{\mathbb{C}}}),$ (5.6)
Proof.
We first split $\Re(\langle\overline{u},{\sf R}(u,h)\rangle)$ into three parts as follows:
$\Re(\langle\overline{u},R(h,u)\rangle)$ $=I_{1}+I_{2}+I_{3}\,$, (5.7)
where
$I_{1}$ $=\Re(\langle\overline{u},\tilde{B}^{*}(u)\rangle)+\Re(\langle\overline{u},D^{1/2}h\otimes D^{1/2}h\rangle)+\Re(\langle\overline{u},D^{1/2}h\otimes D^{1/2}\Upsilon\rangle)\,$,
$I_{2}$ $=-\Re\Big{(}\int_{\mathcal{H}^{+}\cap\left\{0<\|\xi\|_{2}\leq 1\right\}}\big{(}\mathop{}\!\textnormal{e}^{-\langle u,\xi\rangle}-1\big{)}\langle\frac{\mu(\mathop{}\!\textnormal{d}\xi)}{\|\xi\|_{2}^{2}},\overline{u}\rangle\Big{)}\,$,
$I_{3}$ $=-\Re\Big{(}\int_{\mathcal{H}^{+}\cap\left\{\|\xi\|_{2}>1\right\}}\big{(}\mathop{}\!\textnormal{e}^{-\langle u,\xi\rangle}-1\big{)}\langle\frac{\mu(\mathop{}\!\textnormal{d}\xi)}{\|\xi\|_{2}^{2}},\overline{u}\rangle\Big{)}.$
For the first term on the right-hand side of (5.7), it holds
$I_{1}$ $\leq\|B^{*}\|_{\mathcal{L}(\mathcal{H})}\|u\|^{2}_{\mathcal{H}^{\mathbb{C}}}+\|D\|_{\mathcal{L}(H)}\|h\|_{H^{\mathbb{C}}}^{2}\|u\|_{\mathcal{H}^{\mathbb{C}}}+\|D\|_{\mathcal{L}(H)}\|h\|_{H^{\mathbb{C}}}\|\Upsilon\|_{H}\|u\|_{\mathcal{H}^{\mathbb{C}}}$
$\leq g_{1}(1+\|u\|_{\mathcal{H}^{\mathbb{C}}}^{2})(1+\|h\|_{H^{\mathbb{C}}}^{2}),$
where $g_{1}=\|B^{*}\|_{\mathcal{L}(\mathcal{H})}+\|D\|_{\mathcal{L}(H)}+\|D\|_{\mathcal{L}(H)}\|\Upsilon\|_{H}$. Next observe that
$\int_{\mathcal{H}^{+}\cap\left\{\|\xi\|_{2}>1\right\}}\mathop{}\!\textnormal{e}^{-\langle u,\xi\rangle}\frac{\mu(\mathop{}\!\textnormal{d}\xi)}{\|\xi\|_{2}^{2}}$ $\leq_{\mathcal{H}^{+}}\int_{\mathcal{H}^{+}\cap\left\{\|\xi\|_{2}>1\right\}}\mathop{}\!\textnormal{e}^{-\langle\Re(u),\xi\rangle}\frac{\mu(\mathop{}\!\textnormal{d}\xi)}{\|\xi\|_{2}^{2}}$ (5.8)
and
$\int_{\mathcal{H}^{+}\cap\left\{0<\|\xi\|_{2}\leq 1\right\}}(\mathop{}\!\textnormal{e}^{-\langle u,\xi\rangle}-1)\frac{\mu(\mathop{}\!\textnormal{d}\xi)}{\|\xi\|_{2}^{2}}$ $=\int_{\mathcal{H}^{+}\cap\left\{0<\|\xi\|_{2}\leq 1\right\}}\int_{0}^{1}\langle\xi,u\rangle\mathop{}\!\textnormal{e}^{-\langle\xi,u\rangle}\mathop{}\!\textnormal{d}s\frac{\mu(\mathop{}\!\textnormal{d}\xi)}{\|\xi\|_{2}^{2}}$
$\leq_{\mathcal{H}^{+}}\mathop{}\!\textnormal{e}^{\|\Re(u)\|_{2}}\|u\|_{\mathcal{H}^{\mathbb{C}}}\int_{\mathcal{H}^{+}\cap\left\{0<\|\xi\|_{2}\leq 1\right\}}\|\xi\|_{2}^{-1}\mu(\mathop{}\!\textnormal{d}\xi).$ (5.9)
Inequality (5.8) and the monotonicity of $\mathcal{H}^{+}$ imply
$I_{2}$ $\leq\|\int_{\mathcal{H}^{+}\cap\left\{\|\xi\|_{2}>1\right\}}\mathop{}\!\textnormal{e}^{-\langle\Re(u),\xi\rangle}\frac{\mu(\mathop{}\!\textnormal{d}\xi)}{\|\xi\|_{2}^{2}}\|\|u\|_{\mathcal{H}^{\mathbb{C}}}+\|\mu(\mathcal{H}^{+}\cap\left\{\|\xi\|_{2}>1\right\})\|_{2}\|u\|_{\mathcal{H}^{\mathbb{C}}}$
$\leq g_{2}(\Re(u))(1+\|u\|_{\mathcal{H}^{\mathbb{C}}}^{2}),$
where $g_{2}(u)=\|\int_{\mathcal{H}^{+}\cap\left\{\|\xi\|_{2}>1\right\}}\mathop{}\!\textnormal{e}^{-\langle\xi,u\rangle}\frac{\mu(\mathop{}\!\textnormal{d}\xi)}{\|\xi\|_{2}^{2}}\|+\|\mu(\mathcal{H}^{+}\cap\left\{\|\xi\|_{2}>1\right\})\|$. Finally, for the third term in the right hand side of (5.7), it follows from inequality (5.1) and the monotonicity of $\mathcal{H}^{+}$
$I_{3}$ $\leq\|\int_{\mathcal{H}^{+}\cap\left\{0<\|\xi\|_{2}\leq 1\right\}}(\mathop{}\!\textnormal{e}^{-\langle u,\xi\rangle}-1)\frac{\mu(\mathop{}\!\textnormal{d}\xi)}{\|\xi\|_{2}^{2}}\|_{2}\|u\|_{\mathcal{H}^{\mathbb{C}}}.$
$\leq\mathop{}\!\textnormal{e}^{\|\Re(u)\|_{2}}\|\int_{\mathcal{H}^{+}\cap\left\{0<\|\xi\|_{2}\leq 1\right\}}\|\xi\|_{2}^{-1}\mu(\mathop{}\!\textnormal{d}\xi)\|\|u\|_{\mathcal{H}^{\mathbb{C}}}^{2}.$
Hence setting $g_{3}(u)\coloneqq\mathop{}\!\textnormal{e}^{\|u\|_{\mathcal{H}^{\mathcal{C}}}}\|\int_{\mathcal{H}^{+}\cap\left\{0<\|\xi\|_{2}\leq 1\right\}}\|\xi\|_{2}^{-1}\mu(\mathop{}\!\textnormal{d}\xi)\|_{2}$, we find that (5.6) holds for the function $g=g_{1}+g_{2}+g_{3}$, which is locally Lipschitz continuous. ∎
In the following proposition we prove the existence of a unique solution to (2.8a)-(2.8c) on $[0,T]$:
Proposition 5.5.
Let $(b,B,m,\mu)$ be an admissible parameter set satisfying Assumptions Assumption $\mathfrak{A}$ and Assumption $\mathfrak{B}$, let $(\mathcal{A},\operatorname{dom}(\mathcal{A}))$ be the generator of a strongly continuous semigroup, let $D\in\mathcal{L}(H)$ be a positive self-adjoint operator, and let $\Upsilon\in H$. Let $(u_{1},u_{2})\in H^{\mathbb{C}}\times S(\mathcal{O})$. Assume that the extended generalized Riccati equations (2.7a)-(2.7c) have a unique solution on $[0,T]$, for the initial condition $u=(\mathcal{R}(u_{1}),\mathcal{R}(u_{2}))$. Then there exists a unique solution $(\phi(\cdot,u),\psi_{1}(\cdot,u),\psi_{2}(\cdot,u))$ to (2.8a)-(2.8c) on $[0,T]$, with initial value $u=(u_{1},0)$.
Proof.
Let $u=(u_{1},u_{2})\in H^{\mathbb{C}}\times S(\mathcal{O})$ and $T\geq 0$. By standard semigroup theory, we see that the unique mild solution to (2.8c) on $[0,T]$ is given by $\psi_{1}(t,u)=S^{*}(t)u_{1}\in H^{\mathbb{C}}$. For every $t\in[0,T]$ we consider the extension of the map $\mathcal{R}_{u_{1}}$, from the proof of Proposition 5.2, to a function $\mathcal{R}_{u_{1}}(t,\cdot)\colon S(\mathcal{O})\to\mathcal{H}^{\mathbb{C}}$. Then by plugging $\psi_{1}(t,u)$ into (2.7c), we obtain the equation
$\begin{cases}\frac{\partial\psi_{2}}{\partial t}(t,u)&=\mathcal{R}_{u_{1}}(t,\psi_{2}(t,u))\,,\\ \psi_{2}(0,u)&=0.\end{cases}$ (5.10)
The map $\mathcal{R}_{u_{1}}(t,\cdot)$ is locally Lipschitz continuous on $S(\mathcal{O})$, which again by standard arguments implies the existence of a unique solution $\psi_{2}(\cdot,u)$ on some interval $0,T_{\psi_{2}})$ with values in $S(\mathcal{O})\subseteq\mathcal{H}^{\mathbb{C}}$. We want to show that the lifetime $T_{\psi_{2}}$ of $\psi_{2}(\cdot,u_{1},0)$ is always greater or equal than the lifetime $T$ of $q_{2}(\cdot,\Re(u_{1}),0)$. First observe that $\big{(}D^{1/2}S^{*}(t)\Im(u))^{\otimes 2}\leq_{\mathcal{H}^{+}}0$ and
$\int_{\mathcal{H}^{+}\setminus\{0\}}\Re(e^{-\langle\xi,u\rangle}-1)\frac{\mu(\mathop{}\!\textnormal{d}\xi)}{\|\xi\|_{2}^{2}}\leq_{\mathcal{H}^{+}}\int_{\mathcal{H}^{+}\setminus\{0\}}(e^{-\langle\xi,\Re{(u)}\rangle}-1)\frac{\mu(\mathop{}\!\textnormal{d}\xi)}{\|\xi\|_{2}^{2}}.$
This together with the fact that $q_{2}(0,\Re(u_{1}),0)=\Re(\psi_{2}(0,u_{1},0))=0$ imply for all $t<T\wedge T_{\psi_{2}}$
$\frac{\partial q_{2}(t,\Re(u_{1}),0)}{\partial t}-\mathcal{R}_{\Re(u_{1})}\big{(}t,q_{2}(t,\Re(u_{1}),0)\big{)}$
$\qquad=\frac{\partial\Re\psi_{2}(t,u_{1},0)}{\partial t}-\Re\mathcal{R}_{u_{1}}\big{(}t,\psi_{2}(t,u_{1},0)\big{)}$
$\qquad\leq_{\mathcal{H}^{+}}\frac{\partial\Re\psi_{2}(t,u_{1},0)}{\partial t}-\mathcal{R}_{\Re(u_{1})}\big{(}t,\Re\psi_{2}(t,u_{1},0)\big{)}\,.$ (5.11)
Note that $\mathcal{R}_{u}(t,\cdot)$ is quasi-monotone with respect to $\mathcal{H}^{+}$, for $u\in H$ (see [40, Section 6]) for the notion of quasi-monotonicity). Then by Volkmann’s comparison result [40, Satz 2] and (5.1), we have
$q_{2}(t,\Re(u_{1}),0)\leq_{\mathcal{H}^{+}}\Re(\psi_{2}(t,u_{1},0)),\quad\text{for all }t<T\wedge T_{\psi_{2}}.$ (5.12)
We now prove that this implies that $T\leq T_{\psi_{2}}$. First, note that whenever $u\in\mathcal{O}$ and $v\in\mathcal{H}$ are such that $u\leq_{\mathcal{H}^{+}}v$, then $v\in\mathcal{O}$, hence $\Re(\psi_{2}(t,u_{1},0))$ does not approach the boundary of $\mathcal{O}$ before $q_{2}(t,u_{1},0)$ does. Next we show that also $\|\Re\psi_{2}(\cdot,u)\|_{2}$ does not explode before $\|q_{2}(\cdot,u)\|_{2}$. We show this as follows. From (5.8), (5.1), and (5.12), we infer
$\|\Re(\mathcal{R}_{u_{1}}(\psi_{2}(t,u)))\|_{2}$
$\quad\leq\|{B}\|_{\mathcal{L}(\mathcal{H})}\|\Re(\psi_{2}(t,u))\|_{2}+\frac{1}{2}M^{2}\mathop{}\!\textnormal{e}^{2\omega t}\|D\|_{\mathcal{L}(\mathcal{H})}\|u_{1}\|_{H}$
$\qquad+M\mathop{}\!\textnormal{e}^{\omega t}\|D\|_{\mathcal{L}(\mathcal{H})}\|u_{1}\|_{H}\|\Upsilon\|_{H}$
$\qquad+\|\int_{\mathcal{H}^{+}\cap\left\{0<\|\xi\|_{2}\leq 1\right\}}\|\xi\|_{2}^{-1}\mu(\mathop{}\!\textnormal{d}\xi)\|_{2}\mathop{}\!\textnormal{e}^{\|q_{2}(t,\Re(u))\|_{2}}\|\Re{\psi_{2}(t,u)}\|_{2}$
$\qquad+\|\mu(\mathcal{H}^{+}\cap\left\{\|\xi\|_{2}>1\right\})\|_{2}+\|\int_{\mathcal{H}^{+}\cap\left\{\|\xi\|_{2}>1\right\})}\mathop{}\!\textnormal{e}^{-\langle\xi,q_{2}(t,\Re(u))\rangle}\frac{\mu(\mathop{}\!\textnormal{d}\xi)}{\|\xi\|_{2}^{2}}\|_{2},$ (5.13)
where the constants $M\geq 1$ and $\omega\in\mathbb{R}$ are such that $\|S^{*}(t)\|_{\mathcal{L}(H)}\leq M\operatorname{e}^{\omega t}$. Hence by an application of Gronwall’s inequality we conclude that $\|\Re\psi_{2}(t,u)\|_{2}$ is bounded as long as $\|q_{2}(t,u)\|_{2}$ is bounded. Now, by Lemma 5.4, we have
$\Re(\langle\overline{u},R(h,u)\rangle)\leq g(\Re(u))(1+\|u\|_{\mathcal{H}^{\mathbb{C}}}^{2})(1+\|h\|_{H^{\mathbb{C}}}^{2}),$
where $g$ is a locally Lipschitz continuous function on $\mathcal{H}$ and thus
$\frac{\partial}{\partial t}\|\psi_{2}(t,u)\|^{2}$ $=2\Re(\langle\overline{\psi_{2}(t,u)},R(\psi_{1}(t,u),\psi_{2}(t,u))\rangle)$
$\leq g(\Re(\psi_{2}(t,u)))(1+\|\psi_{2}(t,u)\|_{\mathcal{H}^{\mathbb{C}}}^{2})(1+\|\psi_{1}(t,u)\|_{H^{\mathbb{C}}}^{2})\,.$
Again by an application of Gronwall’s inequality and since we already proved that $\Re(\psi_{2}(\cdot,u_{1},0))$ does not explode before $q_{2}(\cdot,\Re(u_{1}),0)$ we conclude that $T_{\psi_{2}}\geq T$, which proves the assertion. ∎
5.2. Analysis of the Riccati equations in the infinite-dimensional Wishart stochastic volatility model
In the following proposition we show the existence of a unique solution to the extended Riccati equations (2.14) up to its maximal lifetime.
Proposition 5.6.
Let $n\in\mathbb{N}$, $Q$, $D\in\mathcal{S}_{1}^{+}(H)$, $\Gamma\in H$ and $\mathfrak{A}\in\mathcal{L}(H)$. Let $(\mathcal{A},\operatorname{dom}(\mathcal{A}))$ be the generator of a strongly continuous semigroup $(S_{t})_{t\geq 0}$ and let $(\mathcal{A}^{*},\operatorname{dom}(\mathcal{A}^{*}))$ be its adjoint. Let $D\in\mathcal{L}(H)$ be a positive self-adjoint operator, and let $\Upsilon\in H$. Then for every $u=(u_{1},u_{2})\in H\times\mathcal{S}_{n}$ there exist a positive real number $T_{\sf q_{2}}$ and a unique solution $(P_{n}(\cdot,u),q_{1}(\cdot,u),q_{2,n}(\cdot,u))$ to (2.14a)-(2.14c) on $0,T_{q_{2}})$.
Proof.
Standard semigroup theory ensures that for any $u_{1}\in H$ and $T>0$ the unique solution to (2.14b) is given by $q_{1,n}(t,u)=S_{t}^{*}u_{1}$ for $t\in[0,T]$. We define $\mathcal{R}_{u_{1}}^{n}(t,\cdot)\colon\mathcal{S}_{n}\to\mathcal{S}_{n}$ by
$\mathcal{R}_{u_{1}}^{n}(t,h)$ $=(\tilde{\mathcal{R}}_{n}(h)-\Pi_{n}\left(\tfrac{1}{2}\big{(}D^{1/2}S_{t}^{*}u_{1}\big{)}^{\otimes 2}-D^{1/2}S_{t}^{*}u_{1}\otimes D^{1/2}\Upsilon\right)\Pi_{n}^{*}\,$, (5.14)
where
$\tilde{R}_{n}(h)=\Pi_{n}\left(\iota_{n}h\iota_{n}^{*}\mathfrak{A}^{*}+\mathfrak{A}\iota_{n}h\iota_{n}^{*}-\frac{1}{2}(\iota_{n}h\iota_{n}^{*}+\iota_{n}h^{T}\iota_{n}^{*})Q(\iota_{n}h\iota_{n}^{*}+\iota_{n}h^{T}\iota_{n}^{*})\right)\Pi_{n}^{*}\,.$
Plugging $q_{1,n}(t,u)$ into (2.14c), we thus obtain the equation
$\begin{cases}\frac{\partial q_{2,n}}{\partial t}(t,u)=\mathcal{R}_{u_{1}}^{n}(t,q_{2,n}(t,u))\,,\\ q_{2,n}(0,u)=u_{2}\,.\end{cases}$
First observe that for $u,v\in\mathcal{S}_{n}$
$\|\tilde{\mathcal{R}}_{n}(u)-\tilde{\mathcal{R}}_{n}(v)\|_{1}$
$\quad\leq 2\|u_{n}-v_{n}\|_{1}\|\mathcal{A}\|_{1}+\frac{1}{2}\|(u_{n}+u_{n}^{T})Q[(u_{n}+u_{n}^{T})-(v_{n}+v_{n}^{T})\|_{1}$
$\qquad+\|[(u_{n}+u_{n}^{T})-(v_{n}+v_{n}^{T})]Q(v_{n}+v_{n}^{T})\|_{1}$
$\quad\leq C(\|u\|_{1}\vee\|v\|_{1})(\|u-v\|_{1})\,$,
where $u_{n}=\iota_{n}u\iota^{*}_{n}$, $v_{n}=\iota_{n}v\iota^{*}_{n}$ and $C$ is a positive constant. It follows that $\mathcal{R}_{u_{1}}(t,\cdot)$ is locally Lipschitz continuous on $\mathcal{S}_{n}$ for every $t\geq 0$ and $u_{1}\in H$. Moreover the function $\tilde{\mathcal{R}}_{n}$ is quasi-monotone with respect to $\mathcal{S}_{n}$ (see [18, Lemma 5.1]). It follows from standard ODE results (see e.g. [36, Chapter 6, Proposition 1.1]) that for every $u_{2}\in\mathcal{S}_{n}$ there exist a $T_{\sf q_{2}}>0$ and a unique solution $q_{2,n}(\cdot,u)$ of (2.14c) on $0,T_{q_{2}})$ with $T_{q_{2}}=\infty$ or
$T_{q_{2}}=\liminf_{p\to\infty}\left\{t\geq 0\colon\|q_{2,n}(t,u)\|_{1}\geq p\right\}.$
By inserting $q_{2,n}(\cdot,u)$ in (2.14a) and observing that $F_{n}$ is continuous on $\mathcal{S}_{n}$, the statement follows. ∎
Appendix A Laguerre polynomials and evaluation of the basis function $f_{k}$
When computing the option prices, one has to calculate the integrals w.r.t. the basis function $f_{k},k=1,\ldots,n$, see for instance, equations (4.9) and (4.17). This requires the evaluations of the basis function $f_{k}$, $k\geq 2$ which themselves are integrals of Laguerre polynomials with weight function $e^{-\frac{1}{2}(\alpha+1)}$, see equation (4.2). We start by introducing the Laguerre polynomials (see, for example, [27, Chapter V] for a good introduction to Laguerre polynomials).
The Laguerre polynomials $L_{n}(x)$ are defined by
$L_{n}(x)=\frac{e^{x}}{n!}\frac{\mathop{}\!\textnormal{d}^{n}}{\mathop{}\!\textnormal{d}x^{n}}\left(x^{n}e^{-x}\right)\,,\qquad n\in\mathbb{N}.$
These polynomials are orthogonal with respect to the weight function $e^{-x}$ on $0,\infty)$, i.e.,
$\int_{0}^{\infty}L_{n}(x)L_{m}(x)e^{-x}\,\mathop{}\!\textnormal{d}x=\begin{cases}0&\text{if }n\neq m,\\ 1&\text{if orthonormalized},\\ \frac{1}{n!}\int_{0}^{\infty}\left[L_{n}(x)\right]^{2}e^{-x}\,\mathop{}\!\textnormal{d}x=\frac{1}{n!}&\text{(optional scaling)}.\end{cases}$
They fulfill the so called three-term recurrence relation
$L_{n+1}(x)=\frac{(2n+1-x)L_{n}(x)-nL_{n-1}(x)}{n+1}\,,\qquad n\geq 2$ (A.1)
with initial conditions
$L_{0}(x)=1,\quad L_{1}(x)=1-x\,.$
The derivative of the standard Laguerre polynomial $L_{n}(x)$ satisfies the identity
$x\frac{\mathop{}\!\textnormal{d}}{\mathop{}\!\textnormal{d}x}L_{n}(x)=nL_{n}(x)-nL_{n-1}(x)\,.$ (A.2)
We make use of these formulas for the Laguerre polynomials to present a fast evaluation approach for $f_{k}$, $k\geq 1$ by iterations. From (A.1), it follows that, for $k\geq 1$
$f_{k+3}(x)$ $=\int_{0}^{x}L_{k+1}(s)e^{-\frac{1}{2}(\alpha+1)s}\mathop{}\!\textnormal{d}s$
$=\frac{2k+1}{k+1}\int_{0}^{x}L_{k}(s)e^{-\frac{1}{2}(\alpha+1)s}\mathop{}\!\textnormal{d}s-\frac{k}{k+1}\int_{0}^{x}L_{k-1}(s)e^{-\frac{1}{2}(\alpha+1)s}\mathop{}\!\textnormal{d}s$
$\qquad-\frac{1}{k+1}\int_{0}^{x}sL_{k}(s)e^{-\frac{1}{2}(\alpha+1)s}\mathop{}\!\textnormal{d}s$
$=\frac{2k+1}{k+1}f_{k+2}(x)-\frac{k}{k+1}f_{k+1}-\frac{1}{k+1}\int_{0}^{x}sL_{k}(s)e^{-\frac{1}{2}(\alpha+1)s}\mathop{}\!\textnormal{d}s.$
Using (A.2), we infer
$\int_{0}^{x}sL_{k}(s)e^{-\frac{1}{2}(\alpha+1)s}\mathop{}\!\textnormal{d}s$
$\quad=-\frac{2}{\alpha+1}sL_{k}(s)e^{-\frac{1}{2}(\alpha+1)s}|_{0}^{x}+\frac{2}{\alpha+1}\int_{0}^{x}[L_{k}(s)+sL_{k}^{\prime}]e^{-\frac{1}{2}(\alpha+1)s}\mathop{}\!\textnormal{d}s$
$\quad=-\frac{2}{\alpha+1}xL_{k}(x)e^{-\frac{1}{2}(\alpha+1)x}+\frac{2}{\alpha+1}f_{k+2}+\frac{2}{\alpha+1}\int_{0}^{x}sL_{k}^{\prime}e^{-\frac{1}{2}(\alpha+1)s}\mathop{}\!\textnormal{d}s$
$\quad=-\frac{2}{\alpha+1}xL_{k}(x)e^{-\frac{1}{2}(\alpha+1)x}+\frac{2}{\alpha+1}f_{k+2}$
$\qquad+\frac{2}{\alpha+1}\int_{0}^{x}[kL_{k}(s)-kL_{k-1}(s)]e^{-\frac{1}{2}(\alpha+1)s}\mathop{}\!\textnormal{d}s$
$\quad=-\frac{2}{\alpha+1}xL_{k}(x)e^{-\frac{1}{2}(\alpha+1)x}+\frac{2}{\alpha+1}(k+1)f_{k+2}-\frac{2}{\alpha+1}kf_{k+1}\,.$
Finally, we have the iteration for the basis functions $f_{k},k\geq 1$
$f_{k+3}(x)$ $=\frac{2k+1}{k+1}f_{k+2}(x)-\frac{k}{k+1}f_{k+1}$
$+\frac{1}{k+1}\frac{2}{\alpha+1}\left[xL_{k}(x)e^{-\frac{1}{2}(\alpha+1)x}-(k+1)f_{k+2}+kf_{k+1}\right]\,.$
Appendix B Numerical method for solving the Riccati equations in the option pricing of Wishart model
In the characteristic function and the option pricing equation, one has to compute the two quantities $\langle X_{t},q_{1}(T-t,u)\rangle,\operatorname{Tr}(Y_{t}q_{2}(T-t,u))$ and P(Tt,u)P(T-t,u). Specially, in implementation we define $t=0$ and $Y_{0}=Q$. Note that $q_{1}(t,u)$ has the same solution as $\psi_{1}(t,u)$ in equation (4.5), and hence $\langle X_{t},q_{1}(T-t,u)\rangle$ can be easily calculated. Moreover, the Riccati equation for $P(t,u)$ is given by
$\frac{\partial P}{\partial t}(t,u)=m\operatorname{Tr}(Qq_{2}(t,u)).$
Note that
$\operatorname{Tr}(Qq_{2}(t,u))$ $=\sum_{n=1}^{\infty}\langle Qq_{2}(t,u)f_{n},f_{n}\rangle$
$=\sum_{n=1}^{\infty}\langle\frac{1}{2}\sum_{k=1}^{\infty}\frac{1}{k^{2}}\langle q_{2}(t,u)f_{n},f_{k}\rangle f_{k},f_{n}\rangle$
$=\frac{1}{2}\sum_{n=1}^{\infty}\frac{1}{n^{2}}\langle q_{2}(t,u)f_{n},f_{n}\rangle.$
Therefore, in order to obtain the characteristic function and the option pricing equation, one has to compute $\langle q_{2}(t,u)f_{n},f_{n}\rangle,n=1,2,\ldots$. Consider the more general case $\langle q_{2}(t,u)f_{i},f_{j}\rangle,i,j=1,2,\ldots$. Recall the Riccati equations for $q_{2}(t,u)$ in the Wishart model are given by
$\frac{\partial q_{2}}{\partial t}(t,u)=Qq_{2}(t,u)+q_{2}(t,u)Q-\frac{1}{2}(D^{\frac{1}{2}}q_{1}(t,u))-\frac{1}{2}(q_{2}(t,u)+q_{2}^{T}(t,u))Q(q_{2}(t,u)+q_{2}^{T}(t,u))$
Discreterize this dynamic, one obtains
$q_{2}(t+\delta,u)$ $=q_{2}(t,u)+\delta(Qq_{2}(t,u)+q_{2}(t,u)Q)-\frac{\delta}{2}(D^{\frac{1}{2}}q_{1}(t,u))^{\otimes 2}$ (B.1)
$-\frac{\delta}{2}\left(q_{2}(t,u)+q_{2}^{T}(t,u))Q(q_{2}(t,u)+q_{2}^{T}(t,u))\right).$
According to equation (B.1), $\langle q_{2}(t,u)f_{i},f_{j}\rangle$ has the iteration
$\langle q_{2}(t+\delta,u)f_{i},f_{j}\rangle-\langle q_{2}(t,u)f_{i},f_{j}\rangle$ (B.2)
$=\delta\langle Qq_{2}(t,u)f_{i},f_{j}\rangle+\delta\langle q_{2}(t,u)Qf_{i},f_{j}\rangle$
$-\frac{\delta}{2}\langle(D^{\frac{1}{2}}q_{1}(t,u))^{\otimes 2}f_{i},f_{j}\rangle-\frac{\delta}{2}\langle(q_{2}(t,u)+q_{2}^{T}(t,u))Q(q_{2}(t,u)+q_{2}^{T}(t,u))f_{i},f_{j}\rangle.$
We will compute the four terms on right hand side of the equation (B.2). First of all, note that
$Qq_{2}(t,u)f_{j}=\frac{1}{2}\sum_{n=1}^{\infty}\frac{1}{n^{2}}\langle q_{2}(t,u)f_{i},f_{n}\rangle f_{n}.$
One obtains
$\langle Qq_{2}(t,u)f_{i},f_{j}\rangle=\frac{1}{2}\langle\sum_{n=1}^{\infty}\frac{1}{n^{2}}\langle q_{2}(t,u)f_{i},f_{n}\rangle f_{n},f_{j}\rangle=\frac{1}{2}\frac{1}{j^{2}}\langle q_{2}(t,u)f_{i},f_{j}\rangle.$ (B.3)
Similarly,
$\langle q_{2}(t,u)Qf_{i},f_{j}\rangle=\frac{1}{2}\frac{1}{i^{2}}\langle q_{2}(t,u)f_{i},f_{j}\rangle.$ (B.4)
For $\langle(D^{\frac{1}{2}}q_{1}(t,u))^{\otimes 2}f_{i},f_{j}\rangle$, note that
$(D^{\frac{1}{2}}q_{1}(t,u))^{\otimes 2}f_{i}$ $=\frac{1}{2}(\nu+\mathrm{i}\lambda)^{2}(\sum_{n=1}^{\infty}\frac{1}{n}f_{n}(t+\vartheta)\langle f_{n},f_{i}\rangle)(\sum_{n=1}^{\infty}\frac{1}{k}f_{k}(t+\vartheta)f_{k})$
$=\frac{1}{2}(\nu+\mathrm{i}\lambda)^{2}\frac{1}{i}f_{i}(t+\vartheta)(\sum_{n=1}^{\infty}\frac{1}{k}f_{k}(t+\vartheta)f_{k}).$
It follows that
$\langle(D^{\frac{1}{2}}q_{1}(t,u))^{\otimes 2}f_{i},f_{j}\rangle=\frac{1}{2}(\nu+\mathrm{i}\lambda)^{2}\frac{1}{ij}f_{i}(t+\vartheta)f_{j}(t+\vartheta).$ (B.5)
Finally, to compute $\langle(q_{2}(t,u)+q_{2}^{T}(t,u))Q(q_{2}(t,u)+q_{2}^{T}(t,u))f_{i},f_{j}\rangle$, we first compute the quantity $(q_{2}(t,u)+q_{2}^{T}(t,u))Q(q_{2}(t,u)+q_{2}^{T}(t,u))f_{i}$ as below:
$(q_{2}(t,u)+q_{2}^{T}(t,u))Q(q_{2}(t,u)+q_{2}^{T}(t,u))f_{i}$
$=\frac{1}{2}(q_{2}(t,u)+q_{2}^{T}(t,u))\sum_{k=1}^{\infty}\frac{1}{k^{2}}\langle q_{2}(t,u)f_{i}+q_{2}^{T}(t,u)f_{i},f_{k}\rangle f_{k}$
$=\frac{1}{2}(q_{2}(t,u)+q_{2}^{T}(t,u))\sum_{k=1}^{\infty}\frac{1}{k^{2}}\left(\langle q_{2}(t,u)f_{i},f_{k}\rangle+\langle q_{2}(t,u)f_{k},f_{i}\rangle\right)f_{k}$
$=\frac{1}{2}\sum_{k=1}^{\infty}\frac{1}{k^{2}}\left(\langle q_{2}(t,u)f_{i},f_{k}\rangle+\langle q_{2}(t,u)f_{k},f_{i}\rangle\right)(q_{2}(t,u)+q_{2}^{T}(t,u))f_{k}$
Consequently
$\langle(q_{2}(t,u)+q_{2}^{T}(t,u))Q(q_{2}(t,u)+q_{2}^{T}(t,u))f_{i},f_{j}\rangle$ (B.6)
$=\frac{1}{2}\sum_{k=1}^{\infty}\frac{1}{k^{2}}\left(\langle q_{2}(t,u)f_{i},f_{k}\rangle+\langle q_{2}(t,u)f_{k},f_{i}\rangle\right)\langle(q_{2}(t,u)+q_{2}^{T}(t,u))f_{k},f_{j}\rangle$
$=\frac{1}{2}\sum_{k=1}^{\infty}\frac{1}{k^{2}}\left(\langle q_{2}(t,u)f_{i},f_{k}\rangle+\langle q_{2}(t,u)f_{k},f_{i}\rangle\right)\left(\langle q_{2}(t,u)f_{k},f_{j}\rangle+\langle q_{2}(t,u)f_{j},f_{k}\rangle\right)$
$=\frac{1}{2}\sum_{k=1}^{\infty}\frac{1}{k^{2}}(\langle q_{2}(t,u)f_{i},f_{k}\rangle\langle q_{2}(t,u)f_{k},f_{j}\rangle+\langle q_{2}(t,u)f_{i},f_{k}\rangle\langle q_{2}(t,u)f_{j},f_{k}\rangle$
$\quad+\langle q_{2}(t,u)f_{k},f_{i}\rangle\langle q_{2}(t,u)f_{k},f_{j}\rangle+\langle q_{2}(t,u)f_{k},f_{i}\rangle\langle q_{2}(t,u)f_{j},f_{k}\rangle)$
Plugging the equations (B.3), (B.4), (B.5) and (B.6) into equation (B.2), one obtains
$\langle q_{2}(t+\delta,u)f_{i},f_{j}\rangle$ $=\langle q_{2}(t,u)f_{i},f_{j}\rangle+\frac{\delta}{2}\frac{1}{j^{2}}\langle q_{2}(t,u)f_{i},f_{j}\rangle+\frac{\delta}{2}\frac{1}{i^{2}}\langle q_{2}(t,u)f_{i},f_{j}\rangle$ (B.7)
$-\frac{\delta}{4}(\nu+\mathrm{i}\lambda)^{2}\frac{1}{ij}f_{i}(t+\vartheta)f_{j}(t+\vartheta)$
$-\frac{\delta}{4}\sum_{k=1}^{\infty}\frac{1}{k^{2}}(\langle q_{2}(t,u)f_{i},f_{k}\rangle\langle q_{2}(t,u)f_{k},f_{j}\rangle+\langle q_{2}(t,u)f_{i},f_{k}\rangle\langle q_{2}(t,u)f_{j},f_{k}\rangle$
$+\langle q_{2}(t,u)f_{k},f_{i}\rangle\langle q_{2}(t,u)f_{k},f_{j}\rangle+\langle q_{2}(t,u)f_{k},f_{i}\rangle\langle q_{2}(t,u)f_{j},f_{k}\rangle).$
In implementation, one has to truncate the infinite sum on the right hand side of equation (B.7) by the chosen dimensionality $n$. Therefore, the iteration in (B.7) is approximated by
$\langle q_{2}(t+\delta,u)f_{i},f_{j}\rangle$ $\approx\langle q_{2}(t,u)f_{i},f_{j}\rangle+\frac{\delta}{2}\frac{1}{j^{2}}\langle q_{2}(t,u)f_{i},f_{j}\rangle+\frac{\delta}{2}\frac{1}{i^{2}}\langle q_{2}(t,u)f_{i},f_{j}\rangle$ (B.8)
$-\frac{\delta}{4}(\nu+\mathrm{i}\lambda)^{2}\frac{1}{ij}f_{i}(t+\vartheta)f_{j}(t+\vartheta)$
$-\frac{\delta}{4}\sum_{k=1}^{n}\frac{1}{k^{2}}(\langle q_{2}(t,u)f_{i},f_{k}\rangle\langle q_{2}(t,u)f_{k},f_{j}\rangle+\langle q_{2}(t,u)f_{i},f_{k}\rangle\langle q_{2}(t,u)f_{j},f_{k}\rangle$
$+\langle q_{2}(t,u)f_{k},f_{i}\rangle\langle q_{2}(t,u)f_{k},f_{j}\rangle+\langle q_{2}(t,u)f_{k},f_{i}\rangle\langle q_{2}(t,u)f_{j},f_{k}\rangle).$
for $i,j=1,\ldots,n$. Denote
$F(t)=\begin{bmatrix}\langle q_{2}(t,u)f_{1},f_{1}\rangle&\langle q_{2}(t,u)f_{1},f_{2}\rangle&\cdots&\langle q_{2}(t,u)f_{1},f_{n}\rangle\\ \langle q_{2}(t,u)f_{2},f_{1}\rangle&\langle q_{2}(t,u)f_{2},f_{2}\rangle&\cdots&\langle q_{2}(t,u)f_{2},f_{n}\rangle\\ \vdots&\vdots&\vdots&\vdots\\ \langle q_{2}(t,u)f_{n},f_{1}\rangle&\langle q_{2}(t,u)f_{n},f_{2}\rangle&\cdots&\langle q_{2}(t,u)f_{n},f_{n}\rangle,\end{bmatrix}$
$B=[\frac{1}{1}f_{1}(t+\vartheta),\frac{1}{2}f_{2}(t+\vartheta),\ldots,\frac{1}{n}f_{n}(t+\vartheta)],$
and
$N=\begin{bmatrix}\frac{1}{1^{2}}&&&\\ &\frac{1}{2^{2}}&&\\ &&\ddots&\\ &&&\frac{1}{n^{2}}\end{bmatrix}$
The Equation (B.8) can be reformed to a matrix form
$F(t+\delta)$ $=F(t)+\frac{\delta}{2}F(t)N+\frac{\delta}{2}NF(t)-\frac{\delta}{4}(\nu+\mathrm{i}\lambda)^{2}B^{T}B$ (B.9)
$-\frac{\delta}{4}(F(t)NF(t)+F(t)NF(t)^{T}+F(t)^{T}NF(t)+F(t)^{T}NF(t)^{T})$
This allows us to iteratively compute $F(t)$ with initial values $F(0)=0$. Therefore, the desired value $\operatorname{Tr}(Qq_{2}(t,u))$ can be computed by obtaining $\langle q_{2}(t,u)f_{k},f_{k}\rangle,k=1,\ldots,n$ from the diagonal of $F(t)$.